### v2

# Sample complexity of optimal transport barycenters with discrete support

Abstract
Computational implementation of optimal transport barycenters for a set of target probability measures requires a form of approximation, a widespread solution being empirical approximation of measures. We provide an $O(\sqrt{N/n})$ statistical generalization bound for the empirical sparse optimal transport barycenters problem, where $N$ is the maximum cardinality of the barycenter (sparse support) and $n$ is the sample size of the target measures empirical approximation. Our analysis includes various optimal transport divergences including Wasserstein, Sinkhorn and Sliced-Wasserstein. We discuss the optimality of our bound as well as the application of our results to specific settings such as K-means, constrained K-means, free and fixed support Wasserstein barycenters.
1 Introduction
Optimal transport barycenters provide a natural way for averaging a set of probability distributions while preserving their underlying geometric structure. They are formally defined as probability measures that minimize the following cost function
$\nu\longmapsto\frac{1}{L}\sum_{\ell=1}^{L}\mathbb{D}\left(\mu^{\ell},\nu\right),$ (1.1)
where $\mathbb{D}$ denotes an optimal transport divergence, and $\left\{\mu^{\ell}\right\}_{\ell=1}^{L}$ is a collection of target probability measures in $\mathbb{R}^{d}$. This problem was first introduced and studied with the 2-Wasserstein metric squared $\mathbb{D}=W_{2}^{2}$ agueh2011barycenters, as a way to generalize McCann’s interpolation mccann1997convexity to more than two measures. Independently and in parallel, the authors of rabin2011wasserstein have proposed a method based on Wasserstein barycenters for texture synthesis and mixing. Then, following fast implementation of $W_{2}$-barycenters (e.g. cuturi2014fast), optimal transport barycenters became a central tool in machine learning and statistics. They have been used for various applications such as sample blending srivastava2015wasp, texture mixing rabin2011wasserstein, multisource domain adaptation montesuma2021wasserstein or music genre recognition montesuma2021wasserstein2. Although initially formulated with $\mathbb{D}=W_{2}^{2}$, other transport divergences have since been considered in (1.1), either to reduce their computational complexity, to induce desirable smoothing, or even to consider a more complex set of probability measures. This has led to the introduction of entropic chizat2023doubly, Bures (for centered Gaussian distributions) chewi2020gradient, sliced bonneel2015sliced and Gromov Wasserstein barycenters brogat2022learning.
Sample averaged approximation and barycenters
There are many situations where the target probability measures in (1.1) may be intractable, either because they are modeled by continuous variables or because they are discretely supported on a very large number of point. To handle these cases from a computational point of view, one requires a discrete approximation. Typical schemes include histogram approximation (e.g. images), sample average approximation (SAA) montesuma2021wasserstein; claici2018stochastic or online stochastic approximation (SA) dvurechenskii2018decentralize; backhoff2024stochastic. This paper focuses on the sample average approximation of target measures, which primarily arises in two main use cases: the statistical setting where the target distributions are only accessible through empirical samples montesuma2021wasserstein and numerical monte-carlo approximation of known densities puccetti2020computation. In both cases, the targets are discrete empirical measures, and the resulting barycenter is therefore itself discrete. Consequently, most algorithms for computing optimal transport barycenters impose a constraint on the support size, similar to optimal quantization of measures pages2004optimal, which consists of approximating a probability measure with a small number of atoms. We include this support constraint in our treatement of the barycenter problem. In other words, we are interested in probability measures $\nu$ minimizing the function (1.1), subject to the constraint $|\operatorname{Supp}(\nu)|\leq N$ for some integer $N\geq 1$, where $\operatorname{Supp}(\nu)$ denotes the support of $\nu$. This problem is referred to as optimal transport barycenters with sparse support in the literature yang2024approximate; borgwardt2021computational. Many modern machine learning problems leverage this barycenter framework, including applications in LLM ai2025resmoe, flow cytometry datasets representation erell2024, graph clustering etienne2025pasco or cross-machine fault diagnosis lee2025barycenter. There exists few exceptions which do not include the support cardinality constraint; for example, staib2017parallel represents barycenters based on estimates of continuous functions by neural networks.
In this article, we provide finite sample statistical guaranties for the sample averaged approximation of optimal transport barycenters under support cardinality constraints. The transport-based divergences $\mathbb{D}$ considered are the Wasserstein distances, the Sinkhorn divergences (entropic regularization), and the Sliced and max-Sliced Wasserstein distances. In other words, we consider a collection of target probability measures $\{\mu^{\ell}\}_{\ell=1}^{L}$ in $\mathbb{R}^{d}$, such that for each $\ell\in[\![1,L]\!]$, we have access to $n$ random variables $X_{1}^{\ell},\ldots,X_{n}^{\ell}\overset{\text{i.i.d}}{\sim}\mu^{\ell}$, and define their associated empirical measures by $\mu_{n}^{\ell}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}^{\ell}}$. The empirical sparse optimal transport barycenters are given by
$\underset{(Y,\pi)\in(\mathbb{R}^{d})^{N}\times\Delta_{N}}{\operatorname{argmin}}\ \frac{1}{L}\sum_{\ell=1}^{L}\mathbb{D}\left(\mu_{n}^{\ell},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)$ (1.2)
where $\mathbb{D}$ denotes any of the transport divergences stated above, $Y:=(y_{1},\ldots,y_{N})$ is a point cloud in $\mathbb{R}^{d}$ and $\Delta_{N}$ is the (N1)(N-1)-probability simplex.
Statistical results for optimal transport barycenters
In carlier2024quantitative authors study the statistical stability of empirical Wasserstein barycenter for general population measures (that is, measures supported on a set of probabilities), which includes the setting of (1.1). The rate of convergence they obtain suffers from the curse of dimensionality; this is inevitable in all generality as the barycenters inherit the sample complexity of optimal transport fournier2015rate. In heinemann2022randomized the authors consider the more specific framework of empirical barycenters of a collection of discrete target measures. Contrarily to carlier2024quantitative where the rate of convergence of empirical barycenter is $O(n^{-1/3d})$, the rate of convergence in this case is $O(C_{d}n^{-1/2})$, where the dimension of the target measures only appears in the constant $C_{d}$, and not exponentially in the sample size. As mentioned before, the fundamental difference in convergence rate behavior between the two frameworks lies in a similar phenomenon observed in optimal transport statistics, namely, that the statistical complexity of optimal transport adapts to the least complex measure hundrieser2024empirical. More precisely, in general, empirical optimal transport has a convergence rate of the form $O(n^{-1/d})$ fournier2015rate, while optimal transport between two probability measures, at least one of which is discrete, has a rate of convergence of $O(n^{-1/2})$ del2024central.
In the case of population barycenters, that is barycenters of a probability measure $\mathbb{Q}$ supported on the space of probability measures, numerous statistical results have been established bigot2020statistical; le2022fast. Although related, these results fall outside the scope of our study, as in their framework, the target measures $\mu^{\ell}$ are independently sampled from $\mathbb{Q}$. Let us note that in the context of entropy regularized Wasserstein barycenters, the sample average approximation was reported to be more efficient than the stochastic approximation approach in terms of computational complexity dvinskikh2020stochastic.
Contribution
We provide convergence bounds in expectation for the empirical sparse optimal transport barycenter functional (1.2) of the form $O(\sqrt{N/n})$, where $N$ denotes the number of support point of the barycenter, and $n$ is the number of samples per empirical target measure. These bounds are obtained for classical, entropy regularized, Sliced and max-Sliced Wasserstein barycenters, bridging the gap between various results in the literature. From this main result, we derive identical rates for the sample complexity of optimal transport barycenters with sparse support. Our proof techniques combine tools from empirical risk minimization theory and novel bounds on dual variables arising in the dual formulations of transport divergences.
Notations
Throughout this article we will denote respectively $\mathcal{M}_{1}(\Omega)$ and $\mathcal{M}_{1}^{N}(\Omega)$, for some $\Omega\subset\mathbb{R}^{d}$, the set of probability measures on $\Omega$ and the set of probability measures supported on at most $N$ points in $\Omega$. We will denote $\mathbb{B}_{R}$ the closed ball of $\mathbb{R}^{d}$ centered in 0 and of radius $R>0$. Additionally, $\|\cdot\|$ is the Euclidean norm on $\mathbb{R}^{d}$, $\Delta_{N}$ the N1N-1 probability simplex and $\mathbb{S}^{d-1}$ the (d1)(d-1)-sphere in $\mathbb{R}^{d}$.
Organization of the paper
We first introduce Wasserstein distances and divergences based on optimal transport, followed by optimal transport barycenters and their sparse approximation in Section 2. We then discuss their computation in Section 3. Our main results on sample complexity of barycenters are presented in Section 4, and the paper ends with a discussion in Section 5. Dual formulations of the divergences considered, together with technical details and proofs, are postponed to the appendices.
2 Introduction to optimal transport and barycenters
2.1 Optimal transport divergences
Optimal transport has become a powerful tool in machine learning due to its ability to compute distances between arbitrary probability measures (we refer the reader to peyre2019computational for a computational point of view, and to villani_opt_old_new for a more theoretical point of view). In mathematical terms, the optimal transport between two measures $\mu,\nu\in\mathcal{M}_{1}(\mathbb{R}^{d})$ consists in finding a coupling which minimizes the average costs of transporting $\mu$ onto $\nu$ :
$\inf_{\gamma\in\Pi(\mu,\nu)}\ \int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}c(x,y)d\gamma(x,y),$ (2.1)
where $c:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ is a lower semi-continuous cost function and $\Pi(\mu,\nu)$ denotes the set of all probability couplings with marginals $\mu$ and $\nu$. When $c:=D^{p}$ is the $p$-exponent of a distance $D$, for $p\geq 1$, (2.1) defines the $p$-exponent of a distance $W_{p}$ in the space of probability measures, called the Wasserstein distance (villani_opt_old_new, Section 6), and is particularly suited as it metrizes the space of probability measures. In our case, we will consider the Euclidean distance on $\mathbb{R}^{d}$, and the Wasserstein distance is given by
$W_{p}^{p}(\mu,\nu)=\inf_{\gamma\in\Pi(\mu,\nu)}\ \int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|x-y\|^{p}d\gamma(x,y).$ (2.2)
In the semi-discrete setting, where one measure is continuous and the other one is discrete, as considered in our sparse barycenter problem, this infinite dimensional optimization problem can be reformulated as a finite-dimensional maximization problem using the Kantorovich dual formulation of optimal transport (see Appendix A). Still, Wasserstein distances lack practicability as in the discrete setting for instance, solving eq. (2.2) involves addressing a potentially high-dimensional linear programming problem. This has motivated the introduction of new optimal transport-based divergences between probability measures.
The entropic regularization of Wasserstein distances, or Sinkhorn divergences cuturi2013sinkhorn; peyre2019computational, consists in adding an entropy regularization term to the optimal transport problem (2.1) and writes between two probability measures $\mu,\nu\in\mathcal{M}_{1}(\mathbb{R}^{d})$ as follows:
$W_{\epsilon,p}^{p}(\mu,\nu)=\inf_{\gamma\in\Pi(\mu,\nu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|x-y\|^{p}d\gamma(x,y)+\epsilon\text{KL}(\gamma|\mu\otimes\nu)$ (2.3)
where $\text{KL}(\gamma|\xi):=\int_{\mathbb{R}^{d^{2}}\times\mathbb{R}^{d^{2}}}(\log(\frac{d\gamma}{d\xi}(x,y))-1)d\gamma(x,y)$ denotes the Kullback-Leibler divergence and $\epsilon>0$ is a regularization parameter. The optimization problem in (2.3) can then be efficiently solved using Sinkhorn algorithm sinkhorn1967concerning; cuturi2013sinkhorn. The dual formulation of this entropy regularized optimal transport problem, studied in genevay2016stochastic with an emphasis on the discrete and semi-discrete settings, will be of particular use in our analysis and is presented in Appendix A. Note that it may be preferable to consider the debiased version of the Sinkhorn divergence genevay2018learning as entropy regularization can produce undesirable smoothing effect when computing sparse barycenters janati2020debiased. It is defined by
$\overline{W}_{\epsilon,p}^{p}(\mu,\nu)=W_{\epsilon,p}^{p}(\mu,\nu)-\frac{1}{2}(W_{\epsilon,p}^{p}(\mu,\mu)+W_{\epsilon,p}^{p}(\nu,\nu)).$ (2.4)
In any case, both classical and regularized optimal transport require solving an optimization problem as soon as the probability distributions lies in an Euclidean space of dimension greater than 1. For the univariate distributions case however, the Wasserstein distance (2.2) has an explicit expression which depends on the quantile function of the measures (ref_villani_topics, Chapter 2.2). By leveraging this analytic form through projections of the $d$-dimensional measures onto lines, rabin2011wasserstein have introduced sliced Wasserstein distances, which were then theoretically studied in kolouri2019generalized. More precisely, for any $\theta\in\mathbb{S}^{d-1}$, we set $P_{\theta}:x\in\mathbb{R}^{d}\mapsto\langle\theta,x\rangle\in\mathbb{R}$ the projection on the line directed by $\theta$. For $\mu,\nu$ probability measures in $\mathbb{R}^{d}$, one can then consider two collections of one-dimensional measures $(P_{\theta\sharp}\mu)_{\theta\in\mathbb{S}^{d-1}}$ and $(P_{\theta\sharp}\nu)_{\theta\in\mathbb{S}^{d-1}}$, where $T_{\sharp}\mu$ denotes the push forward (or image measure) of the measure $\mu$ by the measurable map $T$. For a fixed $\theta$, a sliced distance compares pairs $(P_{\theta\sharp}\mu,P_{\theta\sharp}\nu)$ of one-dimensional measures. If one averages these distances over all directions $\theta\in\mathbb{S}^{d-1}$, then the distance is called Sliced Wasserstein. If one chooses the direction that yields the larger distance between the projected measures, it is called max-Sliced Wasserstein. The Sliced and max-Sliced Wasserstein distance are therefore respectively defined as
$\begin{split}SW_{p}^{p}(\mu,\nu)&=\int_{\mathbb{S}^{d-1}}W_{p}^{p}({P_{\theta}}_{\sharp}\mu,{P_{\theta}}_{\sharp}\nu)d\sigma(\theta)\\ \text{max-}SW_{p}^{p}(\mu,\nu)&=\max_{\theta\in\mathbb{S}^{d-1}}W_{p}^{p}({P_{\theta}}_{\sharp}\mu,{P_{\theta}}_{\sharp}\nu),\end{split}$ (2.5)
where $\sigma$ denotes the uniform probability measure over $\mathbb{S}^{d-1}$. Using the semi-dual formulation of $W_{p}$ one can also express these divergences in the semi-discrete case with a finite dimensional optimization program. These expressions are discussed in Appendix A.
2.2 Optimal transport barycenters with sparse support
An unconstrained barycenter of a collection of target probability measures $\{\mu^{\ell}\}_{\ell=1}^{L}$ with respect to an optimal transport divergence $\mathbb{D}$ is given as a minimizer of (1.1) over $\mathcal{M}_{1}(\mathbb{R}^{d})$. However, for computational reasons discussed in Section 3, this barycenter $\mu^{*}$ may be intractable in practice. For instance, for $\mathbb{D}=W_{p}^{p}$ with $p>1$, and if at least one of the target measures is absolutely continuous with respect to the Lebesgue measure, then the unique barycenter is also absolutely continuous agueh2011barycenters; brizzi2025p, and therefore inherently hard to recover. To address these situations, a common approach is to approximate $\mu^{*}$ by a discrete probability measure by restricting the optimization set in (1.1) to $\mathcal{M}_{1}^{N}(\mathbb{R}^{d})$. This approach is especially relevant when the target measures are available in discretized form (e.g, as empirical measures in an SAA framework), since the resulting barycenter is naturally discrete as well, see Section 5.2. In the literature, this is referred to as optimal transport barycenter problem with sparse support, or equivalently, discretely supported barycenters with a cardinality constraint. Formally, the aim is to find a probability measure $\nu^{*}:=\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}$ where $Y:=(y_{1},\ldots,y_{N})\in(\mathbb{R}^{d})^{N}$ and $\pi\in\Delta_{N}$ , that solves the following optimization problem:
$\underset{(Y,\pi)\in A}{\min}\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right):=\underset{(Y,\pi)\in A}{\min}\frac{1}{L}\sum_{\ell=1}^{L}\mathbb{D}\left(\mu^{\ell},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)$ (2.6)
where A is a closed nonempty subset of $(\mathbb{R}^{d})^{N}\times\Delta_{N}$. When specifying the set of optimization constraint $A$, and the number of measures $L$, one recovers various existing cases in the literature cuturi2014fast; borgwardt2021computational that we summarize in the following Table 1.
AA $L=1$ $L>1$
$(\mathbb{R}^{d})^{N}\times\Delta_{N}$ Optimal quantization Barycenter with sparse support
$(\mathbb{R}^{d})^{N}\times\{\bar{\pi}\}$ Constrained quantization Free support barycenter
$\{\bar{Y}\}\times\Delta_{N}$   Fixed support barycenter
Table 1: Classification of problem (2.6) depending on the constraint set AA and the number of measures $L$.
In this article, we provide statistical complexity bounds for (2.6) with $\mathbb{D}=W_{p}^{p},\>W_{\epsilon,p}^{p},\>\overline{W}_{\epsilon,p}^{p},\>SW_{p}^{p}$ and $\text{max-}SW_{p}^{p}$.
3 Algorithmic aspects of optimal transport barycenters and their sparse approximation
Computing optimal transport barycenters presents two main challenges: the need to discretize the target measures and the inherent computational complexity of optimal transport.
Discretization of target measures
As mentioned before, although the target probability distributions may be continuous in the optimization problem (1.1), an algorithm to compute the barycenter operates on discrete data, as it can only handle finite objects. Generally, continuous target probability measures are therefore approximated by empirical measures, and the barycenter (1.1) is computed either by (deterministic) gradient or fixed-point methods cuturi2014fast, or by stochastic gradient methods li2020continuous. Similarly, for discrete target distributions supported on a large number of atoms, one typically resorts to empirical approximations of the whole distributions and stochastic gradient methods heinemann2022randomized.
Additionally, most practical algorithms implicitly constrain the cardinality of the barycenter, whose number of atoms must be specified as input. Therefore, these algorithms actually solve the constrained problem (2.6) instead of (1.1). This has led to the notion of optimal transport barycenters with sparse support yang2024approximate; borgwardt2021computational. This framework was recently considered in mimouni2025computation, in which the authors propose an algorithm to solve (2.6) for empirical measures. Note that these barycenters are strongly related to optimal quantization and K-means problems (consider $L=1$ in (2.6)), as discussed in Section (5.4).
For (centered) Gaussian target distributions however, a fixed-point algorithm on covariance matrices alvarez2016fixed allows to compute the Gaussian barycenter agueh2011barycenters without discretization. Finally, an alternative to directly handling continuous measures could be to triangulate the supports of the target distributions, as is done in the optimal quantization problem and semi-discrete optimal transport levy2015numerical. However, this approach seems difficult to extend to the barycenter problem since target measures may be distributed over very different supports.
Computational complexity of optimal transport.
Even for discrete probability distributions supported on $N$ points in $\mathbb{R}^{d}$, the optimal transport problem has a prohibitive $O(N^{3}\log(N))$ computational cost peyre2019computational. The Wasserstein barycenter problem naturally inherits this complexity, for a set of $L$ measures altschuler2022wasserstein; borgwardt2021computational. Consequently, the barycenter problem is only tractable for measures supported on a reasonable number of atoms altschuler2021wasserstein.
In practice, Sinkhorn divergences $W_{\epsilon,p}^{p}$ (2.3) are often favored for computing Wasserstein barycenters (1.2), resulting in Sinkhorn barycenters. These divergences offer a significant computational advantage, of order $O(N^{2}/\epsilon)$dvurechensky2018computational, which allows a trade-off between optimal transport accuracy and algorithmic efficiency. Regularization is crucial and can be introduced in two ways : either by using Sinkhorn divergence $\mathbb{D}=W_{\epsilon,p}^{p}$ in the barycenter problem, or by adding an entropic regularization term on the barycenter itself in the optimization problem (1.2). In chizat2023doubly, the author studies both regularization approaches, as well as a combination of the two.
Sliced-Wasserstein distances are also relevant to compute barycenters in the space of probability measures, as introduced in rabin2011wasserstein for an application in texture mixing. The Sliced-Wasserstein barycenters were then studied in bonneel2015sliced. As mentioned in Section 2, Sliced distances rely on the closed-form solution of the one-dimensional Wasserstein distance, and Sliced barycenters consequently inherit this computational efficiency.
4 Sample complexity for sparse optimal transport barycenters
In this section, we present our main results : the generalization error bound followed by the sample complexity of classical, entropic and sliced optimal transport barycenters. These were inspired by (gyorfi2002principles, Section 4.4), which provides the sample complexity of the K-means functional (see Section 5.4).
Theorem 4.1 (Generalization error bound).
Let $\mu^{1},\ldots,\mu^{L}\in\mathcal{M}_{1}(\mathbb{B}_{R})$, for $R>0$. For some integer $n$, let $\mu_{n}^{1},\ldots,\mu_{n}^{L}$ be empirical measures supported over $n$ i.i.d random variables of respective law $\mu^{\ell}$, for all $\ell\in[\![1,L]\!]$. Let $\mathbb{D}=W_{p}^{p},\>W_{\epsilon,p}^{p}$ (for some $\epsilon>0$), $SW_{p}^{p}$ or $\text{max-}SW_{p}^{p}$. Then for any integer $N\geq 1$, and $\operatorname{F}_{\mathbb{D}}$ the function defined in (2.6), we have
$\mathbb{E}\left[\sup_{(Y,\pi)\in\mathbb{B}_{R}^{N}\times\Delta_{N}}\left|\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)-\operatorname{F}_{\mathbb{D}}\left(\mu_{n}^{1},\ldots,\mu_{n}^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)\right|\right]\leq C_{p,R}\sqrt{\frac{C_{d,N}}{n}},$
where the finite constant values $C_{p,R}$ and $C_{d,N}$ depend on the divergence $\mathbb{D}$ and are summarized in the following Table 2.
$\mathbb{D}$ $C_{p,R}$ $C_{d,N}$
$W_{p}^{p}$ $8\sqrt{2}\int_{0}^{2(2R)^{p}}\sqrt{\log\left(2+\frac{32pR^{p}}{\tau}\right)}d\tau$ $N(d+1)$
$SW_{p}^{p}$ $\text{max-}SW_{p}^{p}$ $8\sqrt{2}\int_{0}^{2(2R)^{p}}\sqrt{\log\left(2+\frac{48pR^{p}}{\tau}\right)}d\tau$ $N(d+1)+d$
$W_{\epsilon,p}^{p}$ $8\sqrt{2}\int_{0}^{(4p+1)(2R)^{p}}\sqrt{\log\left(2+\frac{64p(2R)^{p}}{\tau}\right)}d\tau$ $N(d+1)$
Table 2: Explicit constants arising in the generalization error bound for the divergences $\mathbb{D}.$
Theorem 4.1 provides uniform bounds for the empirical barycenter cost function, which is the cost function minimized by most algorithms in practice when computing an optimal transport barycenter (See Section 3).
Remark 4.2 (On the hypothesis of Theorem 4.1).
Note that we consider here empirical measures supported on a set of samples of equal size $n$, but our results also hold for different sample sizes for each target measure by replacing $n$ by the minimum of those individual sample sizes in the upper bound of the generalization error. Finally, although for convenience we consider the barycenter problem (2.6) with uniform weights $\frac{1}{L}$ for each measure $\mu^{\ell}$, Theorem 4.1 holds for any weights $\lambda\in\Delta_{L}$.
A sketch of proof is presented below, with the complete details provided in Appendix C.
Sketch of proof.
We treat here the case $\mathbb{D}=W_{p}^{p}$ for simplicity. Let $(Y,\pi)\in\mathbb{B}_{R}^{N}\times\Delta_{N}$. The first step is to control, for each $\ell=1,\ldots,L$, the deviation
$\left|W_{p}^{p}\!\left(\mu_{n}^{\ell},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)-W_{p}^{p}\!\left(\mu^{\ell},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)\right|.$
This can be achieved by resorting to the dual formulation of $W_{p}^{p}$, in which the dual variables are vectors $w\in\mathbb{R}^{N}$. Lemma D.2 then ensures that there exists functions $f_{Y}^{w}$ such that the above quantity can be upper bounded for each $\ell$ by
$\sup_{w\in\mathbb{R}^{N}}\left|\mathbb{E}_{X\sim\mu_{n}^{\ell}}\!\left[f_{Y}^{w}(X)\right]-\mathbb{E}_{X\sim\mu^{\ell}}\!\left[f_{Y}^{w}(X)\right]\right|.$
At first sight, one cannot apply classical tools from empirical process theory, since the dual variables $w$ are unbounded. Fortunately, Proposition B.1 guarantees that it is enough to restrict the supremum to $w\in[-K,K]^{N}$ for some constant $K$ independent of $Y$ and $\pi$. This boundedness is crucial, as it allows us to invoke standard results from empirical process theory (e.g., (gine2021mathematical, Theorem 3.5.1)) to conclude. ∎
Remark 4.3.
In the specific case where $\mathbb{D}$ is a distance, using the triangular inequality we have that
$|\mathbb{D}(\mu_{n}^{\ell},\nu)-\mathbb{D}(\mu^{\ell},\nu)|\leq\mathbb{D}(\mu_{n}^{\ell},\mu^{\ell}).$ (4.1)
Results on sample complexity of $\mathbb{D}$ could then be used to derive an upper bound on the estimation error in Theorem 4.1. However, for $\mathbb{D}=W_{1}$ for example, the right hand side term of (4.1) is of order $n^{-1/d}$ (chewi2024statistical, Proposition 2.1), which is worse than $n^{-1/2}$ when $d>1$.
Theorem 4.1 directly yields the sample complexity of sparse optimal transport barycenters, as stated in the following corollary.
Corollary 4.4 (Sparse optimal transport barycenters).
Let $\mu^{1},\ldots,\mu^{L}\in\mathcal{M}_{1}(\mathbb{B}_{R})$, for $R>0$. For some integer $n$, let $\mu_{n}^{1},\ldots,\mu_{n}^{L}$ be empirical measures supported over $n$ i.i.d random variables of respective law $\mu^{\ell}$, for all $\ell\in[\![1,L]\!]$. Let also $N\geq 1$ be some integer, $A$ a closed nonempty subset of $\mathbb{B}_{R}^{N}\times\Delta_{N}$, and let $\mathbb{D}=W_{p}^{p},\>W_{\epsilon,p}^{p}$ (for some $\epsilon>0$), $SW_{p}^{p}$ or $\text{max-}SW_{p}^{p}$. We denote by $(Y_{n},\pi_{n})$ an argmin in $A$ of
$(Y,\pi)\mapsto\operatorname{F}_{\mathbb{D}}\left(\mu_{n}^{1},\ldots,\mu_{n}^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)=\frac{1}{L}\sum_{\ell=1}^{L}\mathbb{D}\left(\mu_{n}^{\ell},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right).$
Then we have :
$\begin{split}&\mathbb{E}\left[\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}^{n}\delta_{y_{i}^{n}}\right)-\underset{(Y,\pi)\in A}{\min}\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)\right]\\ &\leq 2C_{p,R}\sqrt{\frac{C_{d,N}}{n}},\end{split}$ (4.2)
where the constant $C_{p,R}$ and $C_{d,N}$ are summarized in the Table 2.
For the specific case of $\mathbb{D}=W_{2}^{2}$ and $L=1$, it was shown in biau2008performance that the estimation error is upper bounded by $O(N/\sqrt{n})$, but with no multiplicative dependence on the dimension $d$. However, this bound is worse by a factor $\sqrt{N}$ compared to our results, for which the rate is of order $O(\sqrt{dN/n})$. The proof of our Corollary 4.4 is obtained leveraging a classical statistical learning theory argument given in the proof below.
Proof.
Suppose $A$ is a closed subset of the compact set $\mathbb{B}_{R}^{N}\times\Delta_{N}$. By continuity of the functionals $(Y,\pi)\mapsto\operatorname{F}_{\mathbb{D}}(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}})$ and $(Y,\pi)\mapsto\operatorname{F}_{\mathbb{D}}(\mu_{n}^{1},\ldots,\mu_{n}^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}})$, they both admit global minimizers in $A$. We denote by $(Y^{*},\pi^{*})$ and $(Y_{n}^{*},\pi_{n}^{*})$ two respective minimizers, and their associated discrete measures $\nu^{*}:=\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}$ and $\nu_{n}^{*}:=\sum_{i=1}^{N}\pi_{i}^{n}\delta_{y_{i}^{n}}$. Then we have
$\begin{split}&\operatorname{F}_{\mathbb{D}}(\mu^{1},\ldots,\mu^{L},\nu_{n}^{*})-\operatorname{F}_{\mathbb{D}}(\mu^{1},\ldots,\mu^{L},\nu^{*})\\ &\leq\operatorname{F}_{\mathbb{D}}(\mu^{1},\ldots,\mu^{L},\nu_{n}^{*})-\operatorname{F}_{\mathbb{D}}(\mu_{n}^{1},\ldots,\mu_{n}^{L},\nu_{n}^{*})+\operatorname{F}_{\mathbb{D}}(\mu_{n}^{1},\ldots,\mu_{n}^{L},\nu^{*})-\operatorname{F}_{\mathbb{D}}(\mu^{1},\ldots,\mu^{L},\nu^{*})\\ &\leq 2\sup_{(Z,\tau)\in A}\left|\operatorname{F}_{\mathbb{D}}\left(\mu_{n}^{1},\ldots,\mu_{n}^{L},\sum_{i=1}^{N}\tau_{i}\delta_{z_{i}}\right)-\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\tau_{i}\delta_{z_{i}}\right)\right|.\end{split}$
Taking the expectation on both sides and applying Theorem 4.1 allows to conclude. ∎
Theorem 4.1 and the sample complexity results of Sinkhorn divergences in genevay2019sample also yield an upper bound for the debiased Sinkhorn barycenters, that is for $\mathbb{D}=\overline{W}_{\epsilon,p}$ defined in (2.4). The proof of the next corollary is also postponed in Appendix C.
Corollary 4.5 (Debiased Sinkhorn barycenters).
Let $\mu^{1},\ldots,\mu^{L}\in\mathcal{M}_{1}(\mathbb{B}_{R})$, for $R>0$. For some integer $n$, let $\mu_{n}^{1},\ldots,\mu_{n}^{L}$ be empirical measures supported over $n$ i.i.d random variables of respective law $\mu^{\ell}$, for all $\ell\in[\![1,L]\!]$. Let also $N\geq 1$ be an integer, $A$ a closed nonempty subset of $\mathbb{B}_{R}^{N}\times\Delta_{N}$ and let $\mathbb{D}=\overline{W}_{\epsilon,p}^{p}$ (for some $\epsilon>0$). We denote by $(Y_{n},\pi_{n})$ an argmin in $A$ of
$(Y,\pi)\mapsto\operatorname{F}_{\mathbb{D}}\left(\mu_{n}^{1},\ldots,\mu_{n}^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right).$
Then we have
$\begin{split}&\mathbb{E}\left[\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}^{n}\delta_{y_{i}^{n}}\right)-\underset{(Y,\pi)\in A}{\min}\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)\right]\\ &\leq C_{p,R}\sqrt{\frac{N(d+1)}{n}}+2K\left(\frac{e^{\frac{{\mathcal{K}}}{\epsilon}}}{\sqrt{n}}\left(1+\frac{1}{\epsilon^{\lfloor d/2\rfloor}}\right)\right),\end{split}$ (4.3)
with $\kappa=2C_{p}R+(2R)^{p}$, where $C_{p}=\sqrt{2}p(2R)^{p-1}$ is the Lipshitz constant of the Euclidean cost at the power $p$. Additionally, $K$ is an unknown constant depending only on $R$ and $p$ and $C_{p,R}$ is the constant associated to $W_{\epsilon,p}^{p}$ in Theorem 4.1.
As the regularization parameter $\epsilon$ goes to zero, the upper bound in (4.3) grows exponentially with the dimension. This arises from the estimation of $W_{\epsilon,p}(\mu^{l},\mu_{n}^{l})$. While the sample complexity of entropic optimal transport remains independent of the dimensionality (even for continuous measures), this no longer holds as $\epsilon\to 0$. In fact, in this limit the curse of dimensionality is implicitly hidden in the regularization parameter $\epsilon$.
5 Discussion of the results
5.1 Sample complexity of optimal transport divergences
Although in general empirical optimal transport suffers from the curse of dimensionality fournier2015rate, it has now been well established that the rate of convergence of the empirical optimal transport divergences between a measure $\mu$ and an empirical measure $\nu_{n}$ distributed over $n$ i.i.d random variables of common law $\nu$, depends on the intrinsic complexity of the measures $\nu$ and $\mu$. If at least one of them is discretely supported, the convergence rate is of order $O(1/\sqrt{n})$ sommerfeld2018inference; genevay2019sample; hundrieser2024empirical; nadjahi2020statistical; lin2021projection, which does not depend exponentially on the dimension of the underlying space $\mathbb{R}^{d}$. In other words, the sample complexity of optimal transport adapts to the measure with lowest complexity hundrieser2024empirical. Our main theorem allows to recover this $\frac{1}{\sqrt{n}}$ rate of convergence for empirical semi-discrete transport. By choosing $L=1$, we get that for any $\nu\in\mathcal{M}_{1}^{N}(\mathbb{B}_{R})$,
$\mathbb{E}|\mathbb{D}(\mu_{n},\nu)-\mathbb{D}(\mu,\nu)|\leq\mathbb{E}\left[\sup_{(Y,\pi)\in\mathbb{B}_{R}^{N}\times\Delta_{N}}\left|\mathbb{D}\left(\mu_{n},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)-\mathbb{D}\left(\mu,\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)\right|\right],$
the upper bounds of Theorem 4.1 also hold for the empirical divergences themselves. We note however that they do not improve upon the literature.
5.2 Barycenters with sparse support
For a sparse barycenter with at most $N$ support points, and empirical target measures supported on $n$ random variables each, the generalization error bounds of our main Theorem 4.1 are of order $\sqrt{N/n}$, which is classical in semi-discrete optimal transport (see (del2024central, Theorem 2.10)). Interestingly, this rate is uniform in the number of measures $L$, in the regularization parameter $\epsilon$ for the Sinkhorn divergence case, and exhibits no curse of dimensionality (i.e no exponential dependency on $d$, the dimension of the ambient space of the target measures).
For a set of discrete target probability measures, a solution of the unconstrained Wasserstein barycenter problem (1.1) for $\mathbb{D}=W_{2}^{2}$ is also a discrete measure. Indeed, assuming that each measure $\mu^{\ell}$ is supported on $n^{\ell}$ points, and setting $N\geq\sum_{\ell=1}^{L}n^{\ell}-L+1$, problems (1.1) and (1.2) coincide: there exists an unconstrained Wasserstein barycenter which turns out to be sparse, supported on a set of maximal cardinality $\sum_{\ell=i}^{L}n^{\ell}-L+1$ (anderes2016discrete, Theorem 2). In this setting, heinemann2022randomized obtains the same rate of convergence $\sqrt{N/\min_{\ell}(n^{\ell})}$ for the unconstrained barycenter, which corresponds to the sparse barycenter considered in this work. This analysis is only limited to large values of $N$.
5.3 Tightness of the result
In this section, we investigate the optimality of the estimation error bound in Corollary 4.4 with respect to the sample size $n$ and the sparsity level $N$.
5.3.1 Lower bound of the estimation error
We aim to establish a lower bound for the estimation error, i.e., the left-hand side term in Corollary 4.4, in a worst-case sense over the choice of target measures. In bartlett2002minimax, the authors consider the quantization problem and a worst case lower bound of the following quantity
$\mathbb{E}\left[W_{2}^{2}\left(\mu,\sum_{i=1}^{N}\bar{\pi}_{i}\delta_{y_{i}^{n}}\right)-\underset{Y\in\mathbb{B}_{R}^{N}}{\min}W_{2}^{2}\left(\mu,\sum_{i=1}^{N}\bar{\pi}_{i}\delta_{y_{i}}\right)\right],$ (5.1)
where $(y_{1}^{n},\ldots,y_{N}^{n})$ denotes a minimizer of $(y_{1},\ldots,y_{N})\mapsto W_{2}^{2}(\mu_{n},\sum_{i=1}^{N}\bar{\pi}_{i}\delta_{y_{i}})$ and $\bar{\pi}$ minimizes the function $\pi\mapsto W_{2}^{2}\left(\mu,\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)$. A direct extension of this result provides a lower bound for the sparse $2$-Wasserstein barycenter problem. The argument is an application of (bartlett2002minimax, Theorem 1), a stronger minimax statement which we only explicit for the empirical barycenter.
Lemma 5.1.
Let $\mathbb{D}=W_{2}^{2}$, the number of measures $L\in\mathbb{N}^{*}$ and the dimension $d\in\mathbb{N}^{*}$. Assume $N\geq 3$ and $n\geq 3435N$ and let $\sum_{i=1}^{N}\pi_{i}^{n}\delta_{y_{i}^{n}}\in\mathcal{M}_{1}^{N}(\mathbb{R}^{d})$ be a random measure depending on $nL$ random variables in $\mathbb{B}_{R}$. Then there exists $\mu^{1},\ldots,\mu^{L}\in\mathcal{M}_{1}(\mathbb{R}^{d})$ such that the estimation error of Corollary 4.4 verifies
$\mathbb{E}\left[\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}^{n}\delta_{y_{i}^{n}}\right)-\min_{(Y,\pi)\in\mathbb{B}_{R}^{N}\times\Delta_{N}}\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)\right]\geq C_{d}\sqrt{\frac{N^{1-\frac{4}{d}}}{n}}$ (5.2)
where $C_{d}$ is an explicit constant which only depends on $d$.
For a fixed $N$, Lemma 5.1 implies that the barycenter functional at a sparse barycenter computed on a dataset of $nL$ observations cannot converge faster than $1/\sqrt{n}$ without additional assumptions. The proof of Lemma 5.1 can be found in Appendix D.2.
We observe the same dependency on $n$ of the estimation error for both the upper bound in (4.2) and the lower bound in (5.2). However the dependency in $N$ differs between these two results. The influence of $N$ on the estimation error depends in fact heavily on the target measures and their intrinsic dimension.
5.3.2 Dependency in $N$ in the estimation error
The lower bound (5.2) of the estimation error is of order $O(N^{-\frac{2}{d}}\sqrt{N/n})$ whereas the upper bound (4.2) is of order $O(\sqrt{N/n})$. Thus, this raises the question of the optimal exponent in $N$, which we investigate in this section by exhibiting a situation where the lower bound is tight.
Let us consider a particular case: $N=n$, $L=1$ and $\pi=(1/n,\ldots,1/n)$, the vector of uniform weights. In this setting, the upper bound (4.2) of Corollary 4.4 does not vanish as $n$ increases. However, we can show that the estimation error is upper bounded by a quantity of order $O(n^{-2/d})$. Indeed, in this setting, the empirical barycenter, i.e. the minimizer of $\mu\mapsto W_{2}^{2}(\mu_{n},\mu)$ over the space $\mathcal{M}_{1}^{N}(\mathbb{R}^{d})$, is clearly $\mu_{n}$ and thus the estimation error writes
$\mathbb{E}\left[W_{2}^{2}(\mu,\mu_{n})-\min_{\nu\in\mathcal{M}_{1}^{n}(\mathbb{B}_{R})}W_{2}^{2}(\mu,\nu)\right].$
This quantity is upper bounded by $\mathbb{E}\left[W_{2}^{2}(\mu,\mu_{n})\right]$, which is of order $O(n^{-2/d})$, see fournier2015rate. This shows that the upper bound derived in Corollary 4.4, which scales in this case as $O\big(1\big)$, is not optimal in that case. As suggested by the above bounding argument, finding a better exponent in $N$ is related to estimating convergence rates for $W_{2}(\mu,\mu_{n})$, a notoriously difficult problem in the statistical optimal transport literature (chewi2024statistical, Section 2). This also indicates that the curse of dimensionality is in fact hidden in the sparsity level $N$, an observation already made in the remark following (del2024central, Theorem 2.10).
5.3.3 Beyond the $W_{2}^{2}$ case
The optimality of the rate of convergence $1/\sqrt{n}$ remains an open question for the other divergences $\mathbb{D}$ which we considered. We expect the same conclusion to hold for $\mathbb{D}=SW_{p}^{p}$ and $\text{max-}SW_{p}^{p}$ since the $1/\sqrt{n}$ rate is also known to be optimal for these divergences xu2022central. Regarding the divergence $\mathbb{D}=W_{\epsilon,p}^{p}$, the empirical Sinkhorn bias benefits from a more favorable $1/n$ convergence rate del2023improved. This observation suggests that a sharper convergence rate might be attainable.
5.4 K-means and constrained K-means
Similarly to the previous section, choosing $\mathbb{D}=W_{2}^{2}$ and $L=1$ in Corollary 4.4 provides error bounds for the celebrated K-means and constrained K-means problems. Let $\mu\in\mathcal{M}_{1}(\mathbb{B}_{R})$ and $X_{1},\ldots,X_{n}$ be $n$ i.i.d samples of law $\mu$. It is well known that the minimization of the K-means functional
$\min_{Y\in(\mathbb{B}_{R})^{N}}\sum_{i=1}^{n}\min_{j=1,\ldots,N}\|X_{i}-y_{j}\|^{2}$ (5.3)
can be recasted as an optimal quantization problem for the empirical measure $\mu_{n}$ distributed on the sample points $X_{1},\ldots,X_{n}$ (see Appendix D.4). In other words, we have
$\min_{Y\in(\mathbb{B}_{R})^{N}}\frac{1}{n}\sum_{i=1}^{n}\min_{j=1,\ldots,N}\|X_{i}-y_{j}\|^{2}=\min_{\nu\in\mathcal{M}_{1}^{N}(\mathbb{B}_{R})}W_{2}^{2}(\mu_{n},\nu).$ (5.4)
Corollary 4.4 then yields a convergence rate of order $O(\sqrt{N/n})$ for the K-means problem, which improves upon the convergence rate $O\left(\sqrt{\frac{N\log(N)}{n}}\right)$ given in (gyorfi2002principles, Section 4.4). Additionally, the Wasserstein functional formulation of the K-means problem provides a complementary approach to the K-means problem, closer to the approximation of measure point of view. It also allows to consider arbitrary cluster size constraints, a variant initially formulated in ng2000note. Formally, the constrained K-means problem consists in fixing a vector of weights $\bar{\pi}\in\Delta_{N}$ and then solving
$\min_{Y\in(\mathbb{R}^{d})^{N}}\ W_{2}^{2}\left(\mu_{n},\sum_{i=1}^{N}\bar{\pi}_{i}\delta_{y_{i}}\right).$ (5.5)
For a known target measure $\mu$ and a uniform vector of weights $\bar{\pi}_{i}=1/N$ for all $i\in[\![1,N]\!]$, we recover the optimal uniform quantization problem which is solved with a variation of the Lloyd algorithm nonassymptotic. In the empirical case $\mu_{n}$, Corollary 4.4 applies to this problem as well.
6 Numerical experiments
In this section, we illustrate  [^1] The code to reproduce the experiments is available in Python at https://github.com/LeoPtls the rate of convergence obtained in Corollary 4.4 for $\mathbb{D}=W_{2}^{2}$ and $W_{\epsilon,2}^{2}$.
Tightness in $n$.
We consider $L=3$ target Gaussian distributions in dimension $d=2$ with random means and random covariance matrices. We compute the $W_{2}^{2}$ and $W_{\epsilon,2}^{2}$ empirical free support barycenter for $N=10$, $N=50$ and for $n$ ranging between $1000$ and $10000$ using the POT library flamary2021pot. The regularization parameter for $W_{\epsilon,2}^{2}$ is set to $\epsilon=1$. Figure 1 displays the evolution of the estimation error in (4.2) as $n$ increases. The observed convergence rate matches the one obtained in Corollary 4.4 for $\mathbb{D}=W_{2}^{2}$ and $\mathbb{D}=W_{\epsilon,2}^{2}$. The presence of hollows in the curve for $\mathbb{D}=W_{\epsilon,p}^{p}$ is due to numerical instabilities.
Figure 1: Estimation error in log-scale as $n$ increases with its associated $95\%$ confidence interval, over $10$ experiments for $\mathbb{D}=W_{2}^{2}$ (top) and $40$ experiments for $\mathbb{D}=W_{\epsilon,2}^{2}$ (bottom). The theoretical bound $n\mapsto 1/\sqrt{n}$ is plotted in orange.
Tightness in $N$.
To illustrate the impact that the sparsity level $N$ has on the estimation error we also consider $L=3$ target Gaussian distributions in dimension $d=6$ with random means and random covariance matrices. We compute the $W_{2}^{2}$ and $W_{\epsilon,2}^{2}$ empirical free support barycenter for fixed sample size $n=500$ and values of $N$ ranging between $10$ and $500$, also using the POT library flamary2021pot. The regularization parameter for $W_{\epsilon,2}^{2}$ is set to $\epsilon=1$. Figure 2 displays the evolution of the estimation error in (4.2). The observed convergence rate lies between the upper bound of Corollary 4.4 and lower bound of Lemma 5.1 for $\mathbb{D}=W_{2}^{2}$, and is close to the upper bound of Corollary 4.4 for $\mathbb{D}=W_{\epsilon,2}^{2}$. This experiment suggests that the dependency in $N$ suggested by the worst case scenario in Lemma 5.1 matches practical observation on mild probability distributions, Gaussian mixtures here.
Figure 2: Estimation error in log-scale as NN increases with its associated $95\%$ confidence interval, over $10$ experiments, and for $n=500$. The theoretical bounds (Corollary 4.4 and Lemma 5.1) are plotted in red.
7 Conclusion
In this paper, we prove that the rate of convergence in expectation of empirical optimal transport barycenters with sparse support is of order $O\left(\sqrt{\frac{N}{n}}\right)$. This result is described for several types of optimal transport divergences, namely the Wasserstein and sliced distances, as well as the Sinkhorn divergence. We show that they all behaved similarly as the number of observation $n$ per target measure grows to infinity. A natural question is whether these rates are min-max optimal. We have showed that this is true for Wasserstein distances $\mathbb{D}=W_{p}^{p}$. However it remains an open question for sliced distances $\mathbb{D}=SW_{p}^{p},\>\text{max-}SW_{p}^{p}$ and Sinkhorn divergences $\mathbb{D}=W_{\epsilon,p}^{p}$. Finally, although our results are limited to the compactly supported case, we believe this is not a significant restriction, as most applications fall within this setting : texture mixing rabin2011wasserstein, constrained clustering ho2017multilevel, shape interpolation solomon2015convolutional.
Appendix A Semi-dual formulation of optimal transport
The Kantorovich dual formulation of the Wasserstein distances $W_{p}$ can be recast as a semi-dual problem using the $c$-transform. Similarly, the Sinkhorn divergences $W_{\epsilon,p}^{p}$ admit a semi-dual formulation based on the smoothed $c$-transform. In the semi-discrete setting (i.e. transport between a continuous and a discrete measure), these formulations turn an infinite-dimensional program (corresponding to the primal formulations (2.1) and (2.3)) into a finite-dimensional one, which is a key point in our proofs. In this section, we present these semi-dual formulations in the semi-discrete setting, with a particular focus on the case where the discrete measure has zero weights or repeated atoms. Throughout this section, we consider a measure $\mu\in\mathcal{M}_{1}(\mathbb{R}^{d})$.
A.1 Semi-dual formulation outside the generalized diagonal
The generalized diagonal plays a central role throughout our proofs, it denotes the set of point clouds with at least two identical components:
$D_{N}=\left\{Y:=(y_{1},\dots y_{N})\in(\mathbb{R}^{d})^{N}\>|\>\exists i\neq j\>\text{such that}\>y_{i}=y_{j}\>\right\}.$ (A.1)
Let us then define $\nu:=\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}$ a discrete measure supported on $Y:=(y_{1},\ldots,y_{N})\in(\mathbb{R}^{d})^{N}\setminus D_{N}$ and strictly positive weights $\pi\in\mathrm{int}(\Delta_{N})$. For a function bounded continuous function $\phi\in\mathcal{C}_{b}(\mathbb{R}^{d})$ , let us also denote $\phi^{c}\in\mathcal{C}_{b}(\mathbb{R}^{d})$ its c-transform, defined in our case for all $x\in\mathbb{R}^{d}$ as $\phi^{c}(x)=\min_{i=1,...,N}(\|x-y_{i}\|^{p}-\phi(y_{i}))$ (book_santambrogio, Section 1.3). Then by Kantorovich’s duality ((villani_opt_old_new, Theorem 5.10)) and $c$-convexity of the optimal potentials we have the following formulation for the $p$-Wasserstein distance between $\mu$ and $\nu$:
$\begin{split}W_{p}^{p}(\mu,\nu):&=\inf_{\gamma\in\Pi(\mu,\nu)}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|x-y\|^{p}d\gamma(x,y)\\ &=\underset{\phi\in\mathbb{R}^{N}}{\max}\int_{\mathbb{R}^{d}}\min_{i=1,\ldots,N}\{\|x-y_{i}\|^{p}-\phi_{i}\}d\mu(x)+\sum_{i=1}^{N}\pi_{i}\phi_{i}.\end{split}$ (A.2)
In (A.2), with a slight abuse of notation, the dual potential $\phi\in\mathcal{C}_{b}(\mathbb{R}^{d})$ is identified with its value on the support of $\nu$, which consists of $N$ distinct points: $\phi_{i}=\phi(y_{i})$, for $i=1,\ldots,N$. For this dual formulation to hold, the measure $\nu$ must be supported outside of the generalized diagonal $D_{N}$, with strictly positive weights; see the discussion in the next Section A.2.
Entropic optimal transport (2.3) between $\mu$ and $\nu$ also admits both a dual and a semi-dual form (see (genevay2016stochastic, Proposition 2.1)), which are respectively given by the following optimization problems:
$W_{\epsilon,p}^{p}(\mu,\nu)=\underset{u,v\in\mathcal{C}(\mathbb{R}^{d})\times\mathbb{R}^{N}}{\max}\int_{\mathbb{R}^{d}}u(x)d\mu(x)+\sum_{i=1}^{N}\pi_{i}v_{i}-\epsilon\sum_{i=1}^{N}\pi_{i}\int_{\mathbb{R}^{d}}\exp\left(\frac{u(x)+v_{i}-\|x-y_{i}\|^{p}}{\epsilon}\right)d\mu(x)$ (A.3)
and
$W_{\epsilon,p}^{p}(\mu,\nu)=\underset{w\in\mathbb{R}^{N}}{\max}-\epsilon\int_{\mathbb{R}^{d}}\log\left(\sum_{i=1}^{N}\pi_{i}\exp\left(\frac{w_{i}-\|x-y_{i}\|^{p}}{\epsilon}\right)\right)d\mu(x)+\sum_{i=1}^{N}\pi_{i}w_{i}-\epsilon.$ (A.4)
These constructions also necessitate that $Y:=(y_{1},\ldots,y_{N})\in(\mathbb{R}^{d})^{N}\setminus D_{N}$ and $\pi\in\mathrm{int}(\Delta_{N})$.
A.2 The case of support points on the generalized diagonal $D_{N}$ and/or with null weights
For a functional of the form $(Y,\pi)\mapsto\mathbb{D}\left(\mu,\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)$, where $\mathbb{D}$ is an optimal transport divergence, it is necessary to assume that the point cloud $Y$ lies outside of the generalized diagonal $D_{N}$, and that the weights are strictly positive in order to apply the duality formulations described above. Indeed, suppose that $y_{i}=y_{j}$, for some $i\neq j$, then the dual formulation of the problem should include the constraint $\phi_{i}=\phi_{j}$, as both $\phi_{i}$ and $\phi_{j}$ represent the evaluation of a continuous potential in the same argument. In other words, we would have $\phi_{i}=\phi(y_{i})=\phi(y_{j})=\phi_{j}$. While this may seem like a minor technicality that can often be overlooked, it cannot be omitted in our analysis. This is particularly relevant in the case of the sliced Wasserstein distance, where two distinct points in $\mathbb{R}^{d}$ may project onto the same point on a line. Similarly, it is important to ensure that the weights of the discrete measure remain in the interior of the probability simplex $\mathrm{int}(\Delta_{N})$.
This technicality has been handled in (portales2024sequential, Section 5.1) with the following proposition.
Proposition A.1.
Let $(Y,\pi)\in(\mathbb{R}^{d})^{N}\times\Delta_{N}$, then there exists $M\in[\![1,N]\!]$ and $(\tilde{Y},\tilde{\pi})\in(\mathbb{R}^{d})^{M}\setminus D_{M}\times\mathrm{int}(\Delta_{M})$ such that for all $\mu\in\mathcal{M}_{1}(\mathbb{R}^{d})$ and any $p\geq 1$:
$W_{p}\left(\mu,\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)=W_{p}\left(\mu,\sum_{i=1}^{M}\tilde{\pi}_{i}\delta_{\tilde{y}_{i}}\right)$
and
$W_{\epsilon,p}\left(\mu,\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)=W_{\epsilon,p}\left(\mu,\sum_{i=1}^{M}\tilde{\pi}_{i}\delta_{\tilde{y}_{i}}\right).$
A.3 Semi-duality formulation for semi-discrete sliced divergences
Starting from the definition of the sliced distances in (2.5), the dual formulation (A.2) can be applied to each one-dimensional distance $W_{p}(P_{\theta\sharp}\mu,P_{\theta\sharp}\nu),\>\theta\in\mathbb{S}^{d-1}$, addressing the ties in the support point of the measures $P_{\theta\sharp}\nu$. Let $(Y,\pi)\in(\mathbb{R}^{d})^{N}\setminus D_{N}\times\mathrm{int}(\Delta_{N})$, then, for each $\theta\in\mathbb{S}^{d-1}$ there exists an integer $M:=M(\theta)\in[\![1,N]\!]$ such that $\langle y_{i}-y_{j},\theta\rangle\neq 0$ and $\pi_{i}>0$ for all $i,j\in[\![1,M(\theta)]\!],\>i\neq j$. Up to a reordering of components, we can write $\pi_{i}$ as $\pi_{i}^{\theta}:=\pi_{i}+\pi_{j}$ whenever $\langle y_{i}-y_{j},\theta\rangle=0$ for all $i,j\in[\![1,N]\!],\>i\neq j$. Following this construction, we obtain a probability measure $\nu^{\theta}:=\sum_{i=1}^{M(\theta)}\pi_{i}^{\theta}\delta_{\langle y_{i},\theta\rangle}$ with no ties in its support points. Then, the Kantorovich duality formulation (A.2) leads to the following expressions of the sliced distances.
$SW_{p}^{p}(\mu,\nu)=\int_{\mathbb{S}^{d-1}}\left\{\underset{w\in\mathbb{R}^{M(\theta)}}{\max}\int_{\mathbb{R}^{d}}\underset{i=1,\ldots,M(\theta)}{\min}\left\{|\langle x-y_{i},\theta\rangle|^{p}-w_{i}\right\}d\mu(x)+\sum_{i=1}^{M(\theta)}\pi_{i}w_{i}\right\}d\sigma(\theta)$ (A.5)
and
$\text{max-}SW_{p}^{p}(\mu,\nu)=\underset{\theta\in\mathbb{S}^{d-1}}{\max}\underset{w\in\mathbb{R}^{M(\theta)}}{\max}\int_{\mathbb{R}^{d}}\underset{i=1,\ldots,M(\theta)}{\min}\left\{|\langle x-y_{i},\theta\rangle|^{p}-w_{i}\right\}d\mu(x)+\sum_{i=1}^{M(\theta)}\pi_{i}w_{i},$ (A.6)
where $\sigma$ denotes the uniform measure over the $d$-sphere $\mathbb{S}^{d-1}$. Note that actually, one can replace $M(\theta)$ by $N$ in the maximization over the variable $w$ in (A.5), without taking into account any dependency on $\theta$. Indeed, we can write
$M(\theta):=\mathrm{card}\left(\left\{i\in[\![1,N]\!]\,:\,\langle y_{i}-y_{j},\theta\rangle\neq 0,\>\forall j\neq i\right\}\right).$
We thus have $M(\theta)<N$ if and only if there exist indices $i\neq j$ such that $\langle y_{i}-y_{j},\theta\rangle=0$ , meaning that $\theta$ lies in
$\underset{i\neq j}{\bigcup_{i,j\in[\![1,N]\!]}}\left\{\theta\in\mathbb{S}^{d-1}\,:\,\langle y_{i}-y_{j},\theta\rangle=0\right\}.$
Since $Y\in(\mathbb{R}^{d})^{N}\setminus D_{N}$ this is a finite union of hyperplanes intersected with the sphere, which is $\sigma$-negligible. It follows that $M(\theta)=N$ for almost every $\theta\in\mathbb{S}^{d-1}$.
Appendix B Boundedness of the dual variables
In this section, we consider a measure $\mu\in\mathcal{M}_{1}(\mathbb{B}_{R}),\>R>0$ and a discrete measure $\nu\in\mathcal{M}_{1}^{N}(\mathbb{B}_{R})$. We prove that the dual variables associated to the optimal transport problem $\mathbb{D}\left(\mu,\nu\right)$ for the Wasserstein distances $\mathbb{D}=W_{p}^{p}$ and the Sinkhorn divergences $\mathbb{D}=W_{\epsilon,p}^{p}$ are bounded.
Proposition B.1.
Let $\mu\in\mathcal{M}_{1}(\mathbb{B}_{R})$ for some $R>0$ and $\mathbb{D}=W_{p}^{p},\>\text{or}\>W_{\epsilon,p}^{p}$. Then for any $(Y,\pi)\in(\mathbb{B}_{R}^{N}\setminus D_{N})\times\mathrm{int}(\Delta_{N})$, we have the following
$W_{p}^{p}\left(\mu,\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)=\underset{\mid w_{i}\mid\leq 2pR^{p}}{\underset{w\in\mathbb{R}^{N}}{\max}}\int_{\mathbb{R}^{d}}\min_{i=1,\ldots,N}\{\|x-y_{i}\|^{p}-w_{i}\}d\mu(x)+\sum_{i=1}^{N}\pi_{i}w_{i},$
and
$W_{\epsilon,p}^{p}\left(\mu,\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)=\underset{\mid w_{i}\mid\leq 4p(2R)^{p}}{\underset{w\in\mathbb{R}^{N}}{\max}}-\epsilon\int_{\mathbb{R}^{d}}\log\left(\sum_{i=1}^{N}\pi_{i}\exp\left(\frac{w_{i}-\|x-y_{i}\|^{p}}{\epsilon}\right)\right)d\mu(x)+\sum_{i=1}^{N}\pi_{i}w_{i}-\epsilon.$
Remark B.2.
In all generality, the dual variables $w\in\mathbb{R}^{N}$ are not bounded since adding a constant $c\in\mathbb{R}$ to $w$ yields another valid solution. Therefore, to fix a dual solution, we set one coordinate of $w$ to 0 and under this constraint the dual variable $w$ remains bounded uniformly in $Y$ and $\pi$.
Proof.
In the following, we prove both cases independently.
Case $\mathbb{D}=W_{p}^{p}$.
Let $(Y,\pi)\in(\mathbb{B}_{R}^{N}\setminus D_{N})\times\mathrm{int}(\Delta_{N})$, then by Kantorovich duality (A.2) between $\mu$ and the discrete measure $\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}$, we have
$W_{p}^{p}\left(\mu,\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)=\max_{w\in\mathbb{R}^{N}}\int_{\mathbb{R}^{d}}\underset{i=1,\ldots,N}{\min}\{\|x-y_{i}\|^{p}-w_{i}\}d\mu(x)+\sum_{i=1}^{N}\pi_{i}w_{i}.$ (B.1)
As explained in Remark (B.2), we may select a choice of dual variables $w\in\mathbb{R}^{N}$ optimal in (B.1) and verifying $w_{N}=0$. From (book_santambrogio, Proposition 5.11) and (book_santambrogio, Remark 1.13) that follows, we may also select a c-convave function $w^{c}$ which verifies for all $i\in[\![1,N]\!]$: $w_{i}=(w^{c}(y_{i}))^{c}$. We then have
$|w_{i}-w_{N}|$ $=|(w^{c}(y_{i}))^{c}-(w^{c}(y_{N}))^{c}|$
$=\left|\min_{x\in\operatorname{Supp}(\mu)}\left\{\|x-y_{i}\|^{p}-w^{c}(x)\right\}-\min_{x\in\operatorname{Supp}(\mu)}\left\{\|x-y_{N}\|^{p}-w^{c}(x)\right\}\right|$
$\leq\max_{x\in\mathbb{B}_{R}}|\|x-y_{i}\|^{p}-\|x-y_{N}\|^{p}|$ (B.2)
$\leq pR^{p-1}\|z_{i}-y_{i}\|$ (B.3)
$\leq 2pR^{p},$
where inequality (B.2) comes from Lemma D.2 and (B.3) from Lemma D.1. Using that $w_{N}=0$, we get the result, namely, the dual variables for the $p$-Wasserstein distances are bounded, and $|w_{i}|\leq 2pR^{p}$.
Case $\mathbb{D}=W_{\epsilon,p}^{p}$.
Let $(Y,\pi)\in(\mathbb{B}_{R}^{N}\setminus D_{N})\times\mathrm{int}(\Delta_{N})$. Let also $u\in\mathcal{C}_{b}(\mathbb{B}_{R})$ be an optimal solution in the dual formulation of the entropic $p$-Wasserstein distance between $\mu$ and $\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}$ given in (A.3). By (genevay2019sample, Proposition 1), $u$ is $\kappa$-Lipshitz, where $\kappa$ is the Lipschitz constant of $(a,b)\in\mathbb{B}_{R}\times\mathbb{B}_{R}\mapsto\|a-b\|^{p}$. Then since $\operatorname{Supp}(\mu)\subset\mathbb{B}_{R}$, we have by Lemma D.1 that $\kappa$ is upper bounded by $p\sqrt{2}(2R)^{p-1}$. Let $w$ be a solution of (A.4) for which we can suppose $w_{N}=0$ (Remark (B.2)). As in genevay2016stochastic, one can show with a first order condition on (A.3) that for all $y$ a coordinate in $Y$:
$w(y)=-\epsilon\log\left(\int_{\mathbb{R}^{d}}\exp\left(\frac{u(x)-\|x-y\|^{p}}{\epsilon}\right)d\mu(x)\right).$
We thus have for any $y_{1},y_{2}$ two distinct points in the point cloud $Y$,
$|w(y_{1})-w(y_{2})|$ $=\left|\epsilon\log\left(\frac{\int_{\mathbb{R}^{d}}\exp\left(\frac{u(x)-\|x-y_{1}\|^{p}}{\epsilon}\right)d\mu(x)}{\int_{\mathbb{R}^{d}}\exp\left(\frac{u(x)-\|x-y_{2}\|^{p}}{\epsilon}\right)d\mu(x)}\right)\right|$
$\leq\left|\epsilon\log\left(\frac{\int_{\mathbb{R}^{d}}\exp\left(\frac{u(\overline{x}_{1})-\|\overline{x}_{1}-y_{1}\|^{p}}{\epsilon}\right)d\mu(x)}{\int_{\mathbb{R}^{d}}\exp\left(\frac{u(\underline{x}_{2})-\|\underline{x}_{2}-y_{2}\|^{p}}{\epsilon}\right)d\mu(x)}\right)\right|$
$=|u(\overline{x}_{1})-\|\overline{x}_{1}-y_{2}\|^{p}-(u(\underline{x}_{2})-\|\underline{x}_{2}-y_{1}\|^{p})|$
$\leq|u(\overline{x}_{1})-u(\underline{x}_{2})|+|\ \|\overline{x}_{1}-y_{2}\|^{p}-\|\underline{x}_{2}-y_{1}\|^{p}|$
$\leq p\sqrt{2}(2R)^{p-1}\|\overline{x}_{1}-\underline{x}_{2}\|+\|\overline{x}_{1}-y_{2}\|^{p}+\|\underline{x}_{2}-y_{1}\|^{p}$ (B.4)
$\leq p\sqrt{2}(2R)^{p}+2(2R)^{p}$
$\leq 4p(2R)^{p},$
where $\overline{x}_{1}$ and $\underline{x}_{2}$ denote respectively a maximizer of $x\mapsto\exp\left(\frac{u(x)-\|x-y_{1}\|^{p}}{\epsilon}\right)$ and a minimizer of $x\mapsto\exp\left(\frac{u(x)-\|x-y_{2}\|^{p}}{\epsilon}\right)$ over $\mathbb{B}_{R}$. Using that $w(y_{N})=0$, we get that for all $i\in[\![1,N]\!]$,
$|w(y_{i})|\leq 4p(2R)^{p}.$
Appendix C Proofs of the main theorems
In this section, we prove the main result in Theorem 4.1, which is the sample complexity of optimal transport barycenters with sparse support. Throughout the section, $\mathbb{P}_{n}^{\ell}$ for $\ell\in[\![1,L]\!]$ denotes the empirical process
$\mathbb{P}_{n}^{\ell}:f\mapsto\mathbb{P}_{n}^{\ell}f:=\frac{1}{n}\sum_{i=1}^{n}f(X_{i}^{\ell})$
for $n$ i.i.d random variables $X_{1}^{\ell},\ldots,X_{n}^{\ell}$ of law $\mu^{\ell}$. For any subset $\mathcal{F}$ of a normed space $(\Theta,\|\cdot\|)$ and $\tau>0$, we denote $\mathcal{N}(\tau,\mathcal{F},\|\cdot\|)$ the $\tau$-covering number of $\mathcal{F}$ with respect to $\|\cdot\|$. This quantity represents the minimum number of balls of radius $\tau$, with respect to the norm $\|\cdot\|$, needed to fully cover $\mathcal{F}$. We refer the reader to (sen2018gentle, Definition 2.2) for more details on the covering number. Finally, we also denote $\|\cdot\|_{\mathbb{L}^{2}(\mu)}$ the $\mathbb{L}^{2}$ norm with respect to the measure $\mu$ and $\overline{\text{B}_{d}(0,R)}$ the $d$-dimensional closed ball of radius $R$.
The proofs for Wasserstein barycenters (Section C.1), Sliced Wasserstein barycenter (Section C.2) and entropy regularized barycenters (Section C.3) all follow the same approach : we first derive the covering number for a class of functions associated to the dual formulation of the divergences $\mathbb{D}$, and then apply tools from empirical process theory to obtain a generalization error bound for the barycenters.
C.1 Wasserstein barycenter
Lemma C.1.
Let
$\mathcal{F}_{p}=\left\{\left\{f_{Y}^{w}:x\in\mathbb{B}_{R}\mapsto\underset{i=1,...,N}{\min}\|x-y_{i}\|^{p}-w_{i}\right\},\>Y\in\mathbb{B}_{R}^{N},\>w\in\overline{\text{B}_{1}(0,2pR^{p})}^{N}\right\},$ (C.1)
then for all $\tau>0$ and $\mu\in\mathcal{M}_{1}(\mathbb{B}_{R})$:
$\mathcal{N}(\tau,\mathcal{F}_{p},\|\cdot\|_{\mathbb{L}^{2}(\mu)})\leq\left(1+\frac{4pR^{p}}{\tau}\right)^{dN}\left(1+\frac{8pR^{p}}{\tau}\right)^{N}$
Moreover $\mathcal{F}_{p,\mathbb{Q}}$, the subset of functions of $\mathcal{F}_{p}$ whose parameters $(Y,w)$ are rationals, is dense in $\mathcal{F}_{p}$.
Proof.
Let $(Y,w),\>(Z,v)\in\mathbb{B}_{R}^{N}\times[-2pR^{p},2pR^{p}]^{N}$. Then for all $x\in\mathbb{B}_{R}$
$|f_{Y}^{w}(x)-f_{Z,v}(x)|$ $=\left|\underset{i=1,...,N}{\min}\|x-y_{i}\|^{p}-w_{i}-\underset{i=1,...,N}{\min}\|x-z_{i}\|^{p}-v_{i}\right|$
$\leq\underset{i=1,...,N}{\max}|\|x-y_{i}\|^{p}-w_{i}-(\|x-z_{i}\|^{p}-v_{i})|$ (C.2)
$\leq\underset{i=1,...,N}{\max}\left\{\left|\|x-y_{i}\|^{p}-\|x-z_{i}\|^{p}\right|+\mid w_{i}-v_{i}\mid\right\}$
$\leq pR^{p-1}\underset{i=1,...,N}{\max}\|y_{i}-z_{i}\|+\underset{i=1,...,N}{\max}|w_{i}-v_{i}|,$ (C.3)
where inequality (C.2) is due to Lemma D.2 and inequality (C.3) is given by Lemma D.1. Let us denote $T=pR^{p-1}$, and let $\tau>0$. We then consider $z^{1},\ldots,z^{n(\tau)}$ a $\frac{\tau}{2T}$-covering of $\mathbb{B}_{R}$ and $v^{1},\ldots v^{m(\tau)}$ a $\frac{\tau}{2}$-covering of $\overline{\text{B}_{1}(0,2pR^{p})}$. For any vector of indices $i=(i_{1},\ldots,i_{N})\in[\![1,n(\tau)]\!]^{N}$ and $k=(k_{1},\ldots,k_{N})\in[\![1,m(\tau)]\!]^{N}$, we define for all $x\in\mathbb{B}_{R}$ the function $f_{i,k}(x)=\underset{j=1,...,N}{\min}\|x-z^{i_{j}}\|^{p}-v^{k_{j}}$. Let $f:=f_{Y}^{w}\in\mathcal{F}_{p}$, then by (C.3) there exists $i,k$ such that
$\begin{split}|f(x)-f_{i,k}(x)|&\leq pR^{p-1}\max_{j=1,...,N}\|y_{j}-z^{i_{j}}\|+\max_{j=1,...,N}|w_{j}-v^{k_{j}}|\\ &\leq T\frac{\tau}{2T}+\frac{\tau}{2}=\tau,\end{split}$
so that $\|f-f_{i,k}\|_{\mathbb{L}^{2}(\mu)}\leq\tau$.
Since there exists $n(\tau)^{N}\times m(\tau)^{N}$ possibilities for such functions $f_{i,k}$, it follows from the covering construction that :
[…]
where we second inequality is due to (massart2007concentration, Lemma 4.14).
Finally, $\mathcal{F}_{p,\mathbb{Q}}$ is dense in $\mathcal{F}_{p}$ by a continuity argument and inequality (C.3). ∎
Theorem C.2.
Let $\mu^{1},\ldots,\mu^{L}\in\mathcal{M}_{1}(\mathbb{B}_{R})$ for some $R>0$. For some integer $n$, let $\mu_{n}^{1},\ldots,\mu_{n}^{L}$ be empirical measures supported over $n$ i.i.d random variables of respective law $\mu^{\ell}$ for all $\ell\in[\![1,L]\!]$, then for $\mathbb{D}=W_{p}^{p}$ and some integer $N\geq 1$, we have
$\begin{split}&\mathbb{E}\left[\sup_{(Y,\pi)\in\mathbb{B}_{R}^{N}\times\Delta_{N}}\left|\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)-\operatorname{F}_{\mathbb{D}}\left(\mu_{n}^{1},\ldots,\mu_{n}^{\ell},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)\right|\right]\\ &\leq C_{p,R}\sqrt{\frac{N(d+1)}{n}},\end{split}$
where $C_{p,R}=8\sqrt{2}\int_{0}^{2(2R)^{p}}\sqrt{\log\left(2+\frac{32pR^{p}}{\tau}\right)}d\tau$.
Proof.
Let $(Y,\pi)\in\mathbb{B}_{R}^{N}\times\Delta_{N}$ and let $M\in[\![1,N]\!]$ and $(\tilde{Y},\tilde{\pi})\in(\mathbb{B}_{R}^{M}\setminus D_{M})\times\mathrm{int}(\Delta_{M})$ be as in Proposition A.1. For all $\ell\in[\![1,L]\!]$, let $X_{1}^{\ell},\ldots,X_{n}^{\ell}$ be $n$ i.i.d random variables of law $\mu^{\ell}$. Using Proposition B.1 and denoting $K:=2pR^{p}$ we therefore have
$\begin{split}&\left|\operatorname{F}_{\mathbb{D}}\left(\mu_{n}^{1},\ldots,\mu_{n}^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)-\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)\right|\\ &=\left|\frac{1}{L}\sum_{\ell=1}^{L}\left[\underset{\mid\tilde{w}_{i}\mid\leq K}{\underset{\tilde{w}\in\mathbb{R}^{M}}{\max}}\frac{1}{n}\sum_{k=1}^{n}\underset{i=1,\ldots,M}{\min}\{\|X_{k}^{\ell}-\tilde{y}_{i}\|^{p}-\tilde{w}_{i}\}+\sum_{i=1}^{M}\tilde{\pi}_{i}\tilde{w}_{i}\right.\right.\\ &\left.\left.-\underset{\mid\tilde{w}_{i}\mid\leq K}{\underset{\tilde{w}\in\mathbb{R}^{M}}{\max}}\int_{\mathbb{R}^{d}}\underset{i=1,\ldots,M}{\min}\{\|x-\tilde{y}_{i}\|^{p}-\tilde{w}_{i}\}d\mu^{\ell}(x)+\sum_{i=1}^{M}\tilde{\pi}_{i}\tilde{w}_{i}\right]\right|\\ &\leq\frac{1}{L}\sum_{\ell=1}^{L}\underset{\mid\tilde{w}_{i}\mid\leq K}{\underset{\tilde{w}\in\mathbb{R}^{M}}{\max}}\left|\frac{1}{n}\sum_{k=1}^{n}\underset{i=1,\ldots,M}{\min}\{\|X_{k}^{\ell}-\tilde{y_{i}}\|^{p}-\tilde{w}_{i}\}-\int_{\mathbb{R}^{d}}\underset{i=1,\ldots,M}{\min}\{\|x-\tilde{y}_{i}\|^{p}-\tilde{w}_{i}\}d\mu^{\ell}(x)\right|\\ &=\frac{1}{L}\sum_{\ell=1}^{L}\underset{w_{i}=w_{j}\>\text{if}\>y_{i}=y_{j}}{\underset{\mid w_{i}\mid\leq K}{\underset{w\in\mathbb{R}^{N}}{\max}}}\left|\frac{1}{n}\sum_{k=1}^{n}\underset{i=1,\ldots,N}{\min}\{\|X_{k}^{\ell}-y_{i}\|^{p}-w_{i}\}-\int_{\mathbb{R}^{d}}\underset{i=1,\ldots,N}{\min}\{\|x-y_{i}\|^{p}-w_{i}\}d\mu^{\ell}(x)\right|.\end{split}$
where the inequality is verified thanks to by Lemma D.2. The last equality is due to the fact that for all $Y\in\mathbb{B}_{R}^{N}$ and $x\in\mathbb{R}^{d}$, we have
$\min_{k=1,\ldots,M}\|x-y_{k}\|^{p}-w_{k}=\min_{k=1,\ldots,N}\|x-y_{k}\|^{p}-w_{k}$
if $w_{i}=w_{j}$ whenever $y_{i}=y_{j}$. Notice that the dependency in the vector of weights $\pi$ has disappeared.
For any $(x,Y,w)\in\mathbb{B}_{R}\times\mathbb{B}_{R}^{N}\times[-K,K]^{N}$, let $f_{Y}^{w}(x)=\underset{i=1,\ldots,N}{\min}\|x-y_{i}\|^{p}-w_{i}$. We then have
$\begin{split}&\underset{(Y,\pi)\in\mathbb{B}_{R}^{N}\times\Delta_{N}}{\sup}\left|\operatorname{F}_{\mathbb{D}}\left(\mu_{n}^{1},\ldots,\mu_{n}^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)-\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)\right|\\ &\leq\frac{1}{L}\sum_{\ell=1}^{L}\underset{(Y,w)\in\mathbb{B}_{R}^{N}\times[-K,K]^{N}}{\sup}\left|\mathbb{P}_{n}^{\ell}\left(f_{Y}^{w}\right)-\mathbb{E}_{X\sim\mu^{\ell}}\left[f_{Y}^{w}(X)\right]\right|\\ &=\frac{1}{L}\sum_{\ell=1}^{L}\underset{f\in\mathcal{F}_{p}}{\sup}|\mathbb{P}_{n}^{\ell}(f)-\mathbb{E}_{X\sim\mu^{\ell}}[f(X)]|\end{split}$
where $\mathcal{F}_{p}$ is defined in (C.1). Notice further that for any $\ell\in[\![1,L]\!]$,
$\begin{split}\sup_{f\in\mathcal{F}_{p}}\mathbb{P}^{\ell}_{n}f^{2}&=\sup_{(Y,w)\in\mathbb{B}_{R}^{N}\times[-K,K]^{N}}\frac{1}{n}\sum_{k=1}^{n}\left(\min_{i=1,...,N}\|X_{k}^{\ell}-y_{i}\|^{p}-w_{i}\right)^{2}\\ &\leq\frac{1}{n}\sum_{k=1}^{n}\left(\sup_{x,y\in\mathbb{B}_{R}}\|x-y\|^{p}+\sup_{w\in[-K,K]}|w|\right)^{2}\\ &\leq((2R)^{p}+2pR^{p})^{2}\leq(2(2R)^{p})^{2},\end{split}$
which gives $\sqrt{\sup_{f\in\mathcal{F}}\mathbb{P}_{n}f^{2}}\leq 2(2R)^{p}$. By Lemma C.1, $\mathcal{F}_{p}$ is dense in $\mathcal{F}_{p,\mathbb{Q}}$, and we can then apply (gine2021mathematical, Theorem 3.5.1) on empirical processes and covering number. Therefore, denoting $\omega:=\underset{f\in\mathcal{F}}{\sup}\mathbb{P}_{n}f^{2}$, we get for any $\ell\in[\![1,L]\!]$ :
$\begin{split}\mathbb{E}\left[\sqrt{n}\underset{f\in\mathcal{F}_{p}}{\sup}|\mathbb{P}_{n}^{\ell}(f)-\mathbb{E}_{X\sim\mu^{\ell}}[f(X)]|\right]&\leq 8\sqrt{2}\mathbb{E}\left[\int_{0}^{\omega}\sqrt{\log\left(2\mathcal{N}(\tau/2,\mathcal{F}_{p},\|\cdot\|_{\mathbb{L}^{2}(\mu_{n}^{\ell})})\right)}d\tau\right]\\ &\leq 8\sqrt{2}\mathbb{E}\left[\int_{0}^{2(2R)^{p}}\sqrt{\log\left(2\mathcal{N}(\tau/2,\mathcal{F}_{p},\|\cdot\|_{\mathbb{L}^{2}(\mu_{n}^{\ell})})\right)}d\tau\right].\end{split}$
From Lemma C.1 applied to the measure $\mu_{n}^{\ell}$, we get
$\begin{split}\int_{0}^{2(2R)^{p}}\sqrt{\log\left(2\mathcal{N}(\tau/2,\mathcal{F}_{p},\|\cdot\|_{\mathbb{L}^{2}(\mu_{n}^{\ell})})\right)}d\tau&\leq\int_{0}^{2(2R)^{p}}\sqrt{\log\left(2\left(1+\frac{8pR^{p}}{\tau}\right)^{dN}\left(1+\frac{16pR^{p}}{\tau}\right)^{N}\right)}d\tau\\ &\leq\int_{0}^{3(2R)^{p}}\sqrt{\log\left(2\left(1+\frac{16pR^{p}}{\tau}\right)^{N(d+1)}\right)}d\tau\\ &\leq\int_{0}^{2(2R)^{p}}\sqrt{\log\left(\left(2+\frac{32pR^{p}}{\tau}\right)^{N(d+1)}\right)}d\tau\\ &=\sqrt{N(d+1)}\int_{0}^{2(2R)^{p}}\sqrt{\log\left(2+\frac{32pR^{p}}{\tau}\right)}d\tau.\end{split}$
Finally, as this last term does not depend on $\ell$, we get
$\mathbb{E}\left[\frac{1}{L}\sum_{\ell=1}^{L}\underset{f\in\mathcal{F}_{p}}{\sup}|\mathbb{P}_{n}^{\ell}(f)-\mathbb{E}_{X\sim\mu^{\ell}}[f(X)]|\right]\leq C_{p,R}\sqrt{\frac{N(d+1)}{n}},$
where $C_{p,R}=8\sqrt{2}\int_{0}^{2(2R)^{p}}\sqrt{\log\left(2+\frac{32pR^{p}}{\tau}\right)}d\tau$. ∎
C.2 Sliced Wasserstein barycenter
Lemma C.3.
Let
$\mathcal{F}_{p}^{s}=\left\{\left\{f_{\theta,Y,w}:x\mapsto\underset{i=1,...,N}{\min}|\langle x-y_{i},\theta\rangle|^{p}-w_{i}\right\},\>Y\in\mathbb{B}_{R}^{N},\>w\in\overline{\text{B}_{1}(0,2pR^{p})}^{N},\theta\in\mathbb{S}^{d-1}\right\},$
then for all $\epsilon>0$ and $\mu\in\mathcal{M}_{1}(\mathbb{B}_{R})$:
$\mathcal{N}(\tau,\mathcal{F}_{p}^{s},\|\cdot\|_{\mathbb{L}^{2}(\mu)})\leq\left(1+\frac{12pR^{p}}{\tau}\right)^{d}\left(1+\frac{6pR^{p}}{\tau}\right)^{dN}\left(\frac{12pR^{p}}{\tau}\right)^{N}.$
Moreover $\mathcal{F}_{p,\mathbb{Q}}^{\text{s}}$, the subset of functions of $\mathcal{F}_{p}^{\text{s}}$ whose parameters $(Y,w,\theta)$ are rationals, is dense in $\mathcal{F}_{p}^{\text{s}}$.
Proof.
Let $K=2pR^{p}$, $Y,Z\in\mathbb{B}_{R}^{N}$, $w,v\in[-K,K]^{N}$ and $\theta,\varphi\in\mathbb{S}^{d-1}$. Then by Lemma (D.1), we have for $x\in\mathbb{B}_{R}$
$\begin{split}\left|\ |\langle x-y_{i},\theta\rangle|^{p}-|\langle x-z_{i},\varphi\rangle|^{p}\right|&\leq pR^{p-1}|\langle x-y_{i},\theta\rangle-\langle x-z_{i},\varphi\rangle|\\ &=pR^{p-1}|\langle x,\theta-\varphi\rangle+\langle z_{i},\varphi-\theta\rangle+\langle z_{i},\theta\rangle-\langle y_{i},\theta\rangle|\\ &\leq pR^{p-1}(|\langle x,\theta-\varphi\rangle|+|\langle z_{i},\varphi-\theta\rangle|+|\langle z_{i}-y_{i},\theta\rangle|)\\ &\leq pR^{p-1}(2R\|\theta-\varphi\|+\|z_{i}-y_{i}\|).\end{split}$
Therefore, thanks to Lemma (D.2), we get
$|f_{\theta,Y,w}(x)-f_{\varphi,Z,v}(x)|$ $\leq\max_{i=1,\ldots,N}\ \left|\ |\langle x-y_{i},\theta\rangle|^{p}-|\langle x-z_{i},\varphi\rangle|^{p}\right|+|w_{i}-v_{i}|$
$\leq\max_{i=1,\ldots,N}\ pR^{p-1}(2R\|\theta-\varphi\|+\|z_{i}-y_{i}\|)+|w_{i}-v_{i}|$
$\leq 2pR^{p}\|\theta-\varphi\|+pR^{p-1}\max_{i=1,\ldots,N}\|z_{i}-y_{i}\|+\max_{i=1,\ldots,N}|w_{i}-v_{i}|.$ (C.4)
Then, using the same argument as in the proof of Lemma C.1, we can upper bound the covering number $\mathcal{N}(\tau,\mathcal{F}_{p}^{s},\|\cdot\|_{\mathbb{L}^{2}(\mu)})$ by the following product of covering numbers:
[…]
which concludes the proof of the upper bound of the covering number. Finally, $\mathcal{F}_{p,\mathbb{Q}}^{s}$ is dense in $\mathcal{F}_{p}^{s}$ due to inequality (C.4). ∎
Theorem C.4.
Let $\mu^{1},\ldots,\mu^{L}\in\mathcal{M}_{1}(\mathbb{B}_{R})$ for some $R>0$. For some integer $n$, let $\mu_{n}^{1},\ldots,\mu_{n}^{L}$ be empirical measures supported over $n$ i.i.d random variables of respective law $\mu^{\ell}$, for all $\ell\in[\![1,L]\!]$. Then for $\mathbb{D}=SW_{p}^{p}$ or $\text{max-}SW_{p}^{p}$ and for some integer $N\geq 1$, we have
$\mathbb{E}\left[\sup_{(Y,\pi)\in\mathbb{B}_{R}^{N}\times\Delta_{N}}\left|\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)-\operatorname{F}_{\mathbb{D}}\left(\mu_{n}^{1},\ldots\mu_{n}^{\ell},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)\right|\right]\leq C_{p,R}\sqrt{\frac{N(d+1)+d}{n}}.$
where $C_{p,R}=8\sqrt{2}\int_{0}^{2(2R)^{p}}\sqrt{\log\left(2+\frac{48pR^{p}}{\tau}\right)}d\tau$.
Proof.
For the sake of brevity, we only prove the result for $\mathbb{D}=SW_{p}^{p}$ as the $\text{max-}SW_{p}^{p}$ case is very similar. Let $M\in[\![1,N]\!]$ and $(\tilde{Y},\tilde{\pi})\in(\mathbb{B}_{R}^{M}\setminus D_{M})\times\mathrm{int}(\Delta_{N})$ be as in Proposition A.1. For all $\ell\in[\![1,L]\!]$, let $X_{1}^{\ell},\ldots,X_{n}^{\ell}$ be i.i.d random variables of law $\mu^{\ell}$. With a slight abuse of notation and following the construction in Section A.3, we define the function $M:\theta\mapsto\operatorname{Card}(\{i\in[\![1,M]\!],\>\langle y_{i}-y_{j},\theta\rangle\neq 0,\forall j\neq i\})$. Then using Proposition B.1 and defining $K:=2pR^{p}$, we have:
$\begin{split}&\left|\operatorname{F}_{\mathbb{D}}\left(\mu_{n}^{1},\ldots,\mu_{n}^{L},\sum_{i=1}^{N}\tilde{\pi}_{i}\delta_{\tilde{y_{i}}}\right)-\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\tilde{\pi}_{i}\delta_{\tilde{y_{i}}}\right)\right|\\ &=\left|\frac{1}{L}\sum_{\ell=1}^{L}\left[\int_{\mathbb{S}^{d-1}}\underset{|w_{i}|\leq K}{\underset{w\in\mathbb{R}^{M(\theta)}}{\max}}\frac{1}{n}\sum_{k=1}^{n}\underset{i=1,\ldots,M(\theta)}{\min}\{|\langle X_{k}^{\ell}-\tilde{y}_{i},\theta\rangle|^{p}-w_{i}\}d\sigma(\theta)+\sum_{i=1}^{M(\theta)}\tilde{\pi}_{i}w_{i}\right.\right.\\ &\left.\left.-\int_{\mathbb{S}^{d-1}}\underset{|w_{i}|\leq K}{\underset{w\in\mathbb{R}^{M(\theta)}}{\max}}\int_{\mathbb{R}^{d}}\underset{i=1,\ldots,M(\theta)}{\min}\{|\langle x-\tilde{y}_{i},\theta\rangle|^{p}-w_{i}\}d\mu^{\ell}(x)d\sigma(\theta)+\sum_{i=1}^{M(\theta)}\tilde{\pi}_{i}w_{i}\right]\right|\\ &\leq\frac{1}{L}\sum_{\ell=1}^{L}\underset{\theta\in\mathbb{S}^{d-1}}{\max}{\underset{|w_{i}|\leq K}{\underset{w\in\mathbb{R}^{M(\theta)}}{\max}}}\left|\frac{1}{n}\sum_{k=1}^{n}\underset{i=1,\ldots,M(\theta)}{\min}\{|\langle X_{k}^{\ell}-y_{i},\theta\rangle|^{p}-w_{i}\}\right.\\ &\left.-\int_{\mathbb{R}^{d}}\underset{i=1,\ldots,M(\theta)}{\min}\{|\langle x-y_{i},\theta\rangle|^{p}-w_{i}\}d\mu^{\ell}(x)\right|.\\ &=\frac{1}{L}\sum_{\ell=1}^{L}\underset{\theta\in\mathbb{S}^{d-1}}{\max}\underset{w_{i}=w_{j}\>\text{if}\>\langle y_{i}-y_{j},\theta\rangle=0}{\underset{|w_{i}|\leq K}{\underset{w\in\mathbb{R}^{N}}{\max}}}\left|\frac{1}{n}\sum_{k=1}^{n}\underset{i=1,\ldots,N}{\min}\{|\langle X_{k}^{\ell}-y_{i},\theta\rangle|^{p}-w_{i}\}\right.\\ &\left.-\int_{\mathbb{R}^{d}}\underset{i=1,\ldots,N}{\min}\{|\langle x-y_{i},\theta\rangle|^{p}-w_{i}\}d\mu^{\ell}(x)\right|,\end{split}$
where the first inequality uses Lemma D.2 for the maximum with respect to $w$, and then upper bound the integral with respect to $\theta$ by its maximum value. The last equality is due to the fact that for all $\theta\in\mathbb{S}^{d-1}$, $Y\in(\mathbb{R}^{d})^{N}$ and $x\in\mathbb{R}^{d}$, we have that
$\min_{k=1,\ldots,M(\theta)}\mid\langle x-y_{k},\theta\rangle\mid^{p}-w_{k}=\min_{k=1,\ldots,N}\mid\langle x-y_{k},\theta\rangle\mid^{p}-w_{k}$
whenever $w_{i}=w_{j}$ if $\langle y_{i}-y_{j},\theta\rangle=0$.
For all $(x,\theta,Y,w)\in\mathbb{B}_{R}\times\mathbb{S}^{d-1}\times\mathbb{B}_{R}^{N}\times[-K,K]^{N}$, let $f_{\theta,Y,w}(x)=\underset{i=1,\ldots,N}{\min}|\langle x-y_{i},\theta\rangle|^{p}-w_{i}$. We then have
$\begin{split}&\underset{(Y,\pi)\in\mathbb{B}_{R}^{N}\times\Delta_{N}}{\sup}\left|\operatorname{F}_{p}\left(\mu_{n}^{1},\ldots,\mu_{n}^{L},\sum_{i=1}^{N}\pi_{i}\delta_{Y_{i}}\right)-\operatorname{F}_{p}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{Y_{i}}\right)\right|\\ &\leq\frac{1}{L}\sum_{\ell=1}^{L}\ \underset{(\theta,Y,w)\in\mathbb{S}^{d-1}\times\mathbb{B}_{R}^{N}\times[-K,K]^{N}}{\sup}\left|\mathbb{P}_{n}^{\ell}\left(f_{\theta,Y,w}\right)-\mathbb{E}_{X\sim\mu^{\ell}}\left[f_{\theta,Y,w}(X)\right]\right|\\ &=\frac{1}{L}\sum_{\ell=1}^{L}\underset{f\in\mathcal{F}_{p}^{\text{s}}}{\sup}|\mathbb{P}_{n}^{\ell}(f)-\mathbb{E}_{X\sim\mu^{\ell}}[f(X)]|,\end{split}$
where $\mathcal{F}_{p}^{\text{s}}$ is defined in Lemma C.3. Moreover, we have for any $\ell\in[\![1,L]\!]$,
$\begin{split}\sup_{f\in\mathcal{F}_{p}^{s}}\mathbb{P}^{\ell}_{n}f^{2}&=\sup_{(\theta,Y,w)\in\mathbb{S}^{d-1}\times\mathbb{B}_{R}^{N}\times[-K,K]^{N}}\frac{1}{n}\sum_{k=1}^{n}\left(\min_{i=1,...,N}|\langle X_{k}^{\ell}-y_{i},\theta\rangle|^{p}-w_{i}\right)^{2}\\ &\leq\frac{1}{n}\sum_{k=1}^{n}\left(\sup_{x,y\in\mathbb{B}_{R},\theta\in\mathbb{S}^{d-1}}\|x-y\|^{p}\|\theta\|^{p}+\sup_{w\in[-K,K]}w\right)^{2}\\ &\leq((2R)^{p}+2pR^{p})^{2}\leq(2(2R)^{p})^{2}.\end{split}$
With the same argument as in the proof of Theorem C.2 this quantity can be upper bounded in expectation using (gine2021mathematical, Theorem 3.5.1):
$\mathbb{E}\left[\sqrt{n}\underset{f\in\mathcal{F}_{p}^{s}}{\sup}\mid\mathbb{P}_{n}^{\ell}(f)-\mathbb{E}_{X\sim\mu^{\ell}}[f(X)]\mid\right]\leq 8\sqrt{2}\mathbb{E}\left[\int_{0}^{2(2R)^{p}}\sqrt{\log\left(2\mathcal{N}(\tau/2,\mathcal{F}_{p}^{\text{s}},\|\cdot\|_{\mathbb{L}^{2}(\mu_{n}^{\ell})})\right)}d\tau\right].$
From Lemma C.3 we thus have
[…]
Finally, as this last term does not depend on $\ell$, we get
$\mathbb{E}\left[\frac{1}{L}\sum_{\ell=1}^{L}\underset{f\in\mathcal{F}_{p}^{\text{s}}}{\sup}|\mathbb{P}_{n}^{\ell}(f)-\mathbb{E}_{X\sim\mu^{\ell}}[f(X)]|\right]\leq C_{p,R}\sqrt{\frac{N(d+1)+d}{n}},$
where $C_{p,R}=\int_{0}^{2(2R)^{p}}\sqrt{\log\left(2+\frac{48pR^{p}}{\tau}\right)}d\tau$.
C.3 Entropy regularized barycenter
Lemma C.5.
Let
$\mathcal{F}_{p}^{\epsilon}=\left\{\left\{g_{Y,\pi,w}^{\epsilon}:x\in\mathbb{B}_{R}\mapsto-\epsilon\log\left(\sum_{i=1}^{N}\pi_{i}\exp\left(\frac{w_{i}-\|x-y_{i}\|^{p}}{\epsilon}\right)\right)\right\}\right$.
$\left.\middle|Y\in\mathbb{B}_{R}^{N},\;\pi\in\Delta_{N},\;w\in\overline{\text{B}_{1}\left(0,4p(2R)^{p}\right)}^{N},\right\}$
then for all $\tau>0$ and $\mu\in\mathcal{M}_{1}(\mathbb{B}_{R})$, we have
$\mathcal{N}(\tau,\mathcal{F}_{p}^{\epsilon},\|\cdot\|_{\mathbb{L}^{2}(\mu)})\leq\left(1+\frac{4pR^{p}}{\tau}\right)^{dN}\left(\frac{16p(2R)^{p}}{\tau}\right)^{N}.\\$
Moreover $\mathcal{F}_{p,\mathbb{Q}}^{\epsilon}$, the subset of functions of $\mathcal{F}_{p}^{\epsilon}$ whose parameters $(Y,\pi,w)$ are rationals, is dense in $\mathcal{F}_{p}^{\epsilon}$.
Proof.
Let $(Y,\pi,w),\>(Z,\gamma,v)\in\mathbb{B}_{R}^{N}\times\Delta_{N}\times[-4p(2R)^{p},4p(2R)^{p}]^{N}$. Then for all $x\in\mathbb{B}_{R}$, we have
$\begin{split}|g_{Y,\pi,w}^{\epsilon}(x)-g_{Z,\gamma,v}^{\epsilon}(x)|&=\epsilon\left|\log\left(\frac{\sum_{i=1}^{N}\pi_{i}\exp\left(\frac{w_{i}-\|x-y_{i}\|^{p}}{\epsilon}\right)}{\sum_{i=1}^{N}\gamma_{i}\exp\left(\frac{v_{i}-\|x-z_{i}\|^{p}}{\epsilon}\right)}\right)\right|\\ &\leq\epsilon\left|\log\left(\frac{\max_{i=1,\ldots,N}\exp\left(\frac{w_{i}-\|x-y_{i}\|^{p}}{\epsilon}\right)\sum_{i=1}^{N}\pi_{i}}{\min_{i=1,\ldots,N}\exp\left(\frac{v_{i}-\|x-z_{i}\|^{p}}{\epsilon}\right)\sum_{i=1}^{N}\gamma_{i}}\right)\right|\\ &=\left|\max_{i=1,\ldots,N}w_{i}-\|x-y_{i}\|^{p}-\min_{i=1,\ldots,N}v_{i}-\|x-z_{i}\|^{p}\right|\\ &\leq|w_{\overline{i}}-v_{\underline{i}}|+|\|x-z_{\underline{i}}\|^{p}-\|x-y_{\overline{i}}\|^{p}|\end{split}$
where $\overline{i}$ (resp. $\underline{i}$) is an index for which the maximum (resp. minimum) of $i\mapsto w_{i}-\|x-y_{i}\|^{p}$ (resp. $i\mapsto v_{i}-\|x-z_{i}\|^{p}$) is attained. By Lemma D.1, we thus have
$|g_{Y,\pi,w}(x)-g_{Z,\gamma,v}(x)|\leq|w_{\overline{i}}-v_{\underline{i}}|+pR^{p-1}\|y_{\overline{i}}-z_{\underline{i}}\|.$ (C.5)
Notice that the right-hand side of the above inequality depends on neither $\pi$ nor $\epsilon$. Let $T=pR^{p-1}$, with the same proof argument of Theorem C.2, and by boundedness of the entropic dual variables (Proposition B.1) we thus get
[…]
which concludes the proof of the upper bound of the covering number. Finally, $\mathcal{F}_{p,\mathbb{Q}}^{\epsilon}$ is dense in $\mathcal{F}_{p}^{\epsilon}$ due to inequality (C.5). ∎
Theorem C.6.
Let $\mu^{1},\ldots,\mu^{L}\in\mathcal{M}_{1}(\mathbb{B}_{R})$ for some $R>0$. For some integer $n$, let $\mu_{n}^{1},\ldots,\mu_{n}^{L}$ be empirical measures supported over $n$ i.i.d random variables of respective law $\mu^{\ell}$, for all $\ell\in[\![1,L]\!]$. Let $\mathbb{D}=W_{\epsilon,p}^{p}$ with $\epsilon>0$, then for some integer $N\geq 1$:
$\mathbb{E}\left[\sup_{(Y,\pi)\in\mathbb{B}_{R}^{N}\times\Delta_{N}}\left|\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)-\operatorname{F}_{\mathbb{D}}\left(\mu_{n}^{1},\ldots\mu_{n}^{\ell},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)\right|\right]\leq C_{p,R}\sqrt{\frac{N(d+1)}{n}}$
where $C_{p,R}=8\sqrt{2}\int_{0}^{(4p+1)(2R)^{p}}\sqrt{\log\left(2+\frac{64p(2R)^{p}}{\tau}\right)}d\tau$.
Proof.
We fix $\epsilon>0$. Let $(Y,\pi)\in\mathbb{B}_{R}^{N}\times\Delta_{N}$ and let $M\in[\![1,N]\!]$ and $(\tilde{Y},\tilde{\pi})\in(\mathbb{B}_{R}^{M}\setminus D_{M})\times\mathrm{int}(\Delta_{M})$ be as in Proposition A.1. For any triplet $(Y,\pi,w)\in\mathbb{B}_{R}^{N}\times\Delta_{N}\times\mathbb{R}^{N}$ and $x\in\mathbb{B}_{R}$, we define the function
$g_{Y,\pi,w}^{\epsilon}(x):=-\epsilon\log\left(\sum_{j=1}^{N}\pi_{j}\exp\left(\frac{w_{j}-\|x-y_{j}\|^{p}}{\epsilon}\right)\right).$
For all $\ell\in[\![1,L]\!]$, let $X_{1}^{\ell},\ldots,X_{n}^{\ell}$ be i.i.d random variables of law $\mu^{\ell}$. Using Proposition B.1 and the same argument as in the proof of Theorem C.2, we get for $K:=4p(2R)^{p}$
$\begin{split}&\underset{(Y,\pi)\in\mathbb{B}_{R}^{N}\times\Delta_{N}}{\sup}\left|\operatorname{F}_{\mathbb{D}}\left(\mu_{n}^{1},\ldots,\mu_{n}^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)-\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)\right|\\ &\leq\frac{1}{L}\sum_{\ell=1}^{L}\ \underset{(Y,\pi,w)\in\mathbb{B}_{R}^{N}\times\Delta_{N}\times[-K,K]^{N}}{\sup}|\mathbb{P}_{n}^{\ell}g_{Y,\pi,w}^{\epsilon}-\mathbb{E}_{X\sim\mu^{\ell}}\left[g_{Y,\pi,w}^{\epsilon}(X)\right]|\\ &=\frac{1}{L}\sum_{\ell=1}^{L}\sup_{g\in\mathcal{F}_{p}^{\epsilon}}|\mathbb{P}_{n}^{\ell}(g)-\mathbb{E}_{X\sim\mu^{\ell}}[g(X)]|\end{split}$
where $\mathcal{F}_{p}^{\epsilon}$ is defined in Lemma C.5 . Moreover, for all $w\in\overline{\text{B}_{1}(0,4p(2R)^{p})}^{N}$, we have
$-\frac{4p(2R)^{p}+(2R)^{p}}{\epsilon}\leq\frac{w_{j}-\|x-y_{j}\|^{p}}{\epsilon}\leq\frac{4p(2R)^{p}}{\epsilon}$
$\iff-\frac{(4p+1)(2R)^{p}}{\epsilon}\leq\log\left(\sum_{j=1}^{N}\pi_{j}\exp\left(\frac{w_{j}-\|x-y_{j}\|^{p}}{\epsilon}\right)\right)\leq\frac{4p(2R)^{p}}{\epsilon}$
$\iff-4p(2R)^{p}\leq g_{Y,\pi,w}^{\epsilon}(x)\leq(4p+1)(2R)^{p}.$ (C.6)
Then, by Lemma C.5, the bound obtained in (C.6) and with the same argument as in Theorem C.2, we obtain by (gine2021mathematical, Theorem 3.5.1) the following bound
[…]
so that finally
$\mathbb{E}\left[\frac{1}{L}\sum_{\ell=1}^{L}\ \underset{g\in\mathcal{F}_{p}^{\epsilon}}{\sup}|\mathbb{P}_{n}^{\ell}(g)-\mathbb{E}_{X\sim\mu^{\ell}}[g(X)]|\right]\leq C_{p,R}\sqrt{\frac{N(d+1)}{n}},$
where $C_{p,R}=8\sqrt{2}\int_{0}^{(4p+1)(2R)^{p}}\sqrt{\log\left(2+\frac{64p(2R)^{p}}{\tau}\right)}d\tau$. ∎
In the following, we prove the result for debiased Sinkhorn barycenters.
Corollary C.7.
Let $\mu^{1},\ldots,\mu^{L}\in\mathcal{M}_{1}(\mathbb{B}_{R})$, for $R>0$. For some integer $n$, let $\mu_{n}^{1},\ldots,\mu_{n}^{L}$ be empirical measures supported over $n$ i.i.d random variables of respective law $\mu^{\ell}$, for all $\ell\in[\![1,L]\!]$. Let also $N\geq 1$ be an integer, $A$ a closed nonempty subset of $\mathbb{B}_{R}^{N}\times\Delta_{N}$ and let $\mathbb{D}=\overline{W}_{\epsilon,p}^{p}$ (for some $\epsilon>0$). We denote by $(Y^{n},\pi^{n})$ an argmin in $A$ of
$(Y,\pi)\mapsto\operatorname{F}_{\mathbb{D}}\left(\mu_{n}^{1},\ldots,\mu_{n}^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right).$
Then we have
$\begin{split}&\mathbb{E}\left[\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}^{n}\delta_{y_{i}^{n}}\right)-\underset{(Y,\pi)\in A}{\min}\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)\right]\\ &\leq 2C_{p,R}\sqrt{\frac{N(d+1)}{n}}+2K_{p,R}\left(\frac{e^{\frac{{\mathcal{K}}}{\epsilon}}}{\sqrt{n}}\left(1+\frac{1}{\epsilon^{\lfloor d/2\rfloor}}\right)\right),\end{split}$
with $\kappa=2C_{p}R+(2R)^{p}$, where $C_{p}=\sqrt{2}p(2R)^{p-1}$ is the Lipshitz constant of the Euclidean cost at the power $p$. Additionally, $K_{p,R}$ is an unknown constant depending only on $R$ and $p$ and $C_{p,R}$ is the constant associated to $W_{\epsilon,p}^{p}$ in Theorem 4.1.
Proof.
Let $(Y^{*},\pi^{*})$ and $(Y^{n},\pi^{n})$ be respective minimizers of $(Y,\pi)\mapsto\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)$ and $(Y,\pi)\mapsto\operatorname{F}_{\mathbb{D}}\left(\mu_{n}^{1},\ldots,\mu_{n}^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)$. We then denote $\nu^{*}=\sum_{i=1}^{N}\pi_{i}^{*}\delta_{y_{i}^{*}}$ and $\nu^{n}=\sum_{i=1}^{N}\pi_{i}^{n}\delta_{y_{i}^{n}}$. Note that the minimizers exist by continuity of the functional on the compact set $A$. Notice now that
$\begin{split}&\operatorname{F}_{\mathbb{D}}(\mu^{1},\ldots,\mu^{L},\nu^{n})-\operatorname{F}_{\mathbb{D}}(\mu^{1},\ldots,\mu^{L},\nu^{*})\\ &\leq\operatorname{F}_{\mathbb{D}}(\mu^{1},\ldots,\mu^{L},\nu^{n})-\operatorname{F}_{\mathbb{D}}(\mu_{n}^{1},\ldots,\mu_{n}^{L},\nu^{n})+\operatorname{F}_{\mathbb{D}}(\mu_{n}^{1},\ldots,\mu_{n}^{L},\nu^{*})-\operatorname{F}_{\mathbb{D}}(\mu^{1},\ldots,\mu^{L},\nu^{*})\\ &\leq 2\underset{(Y,\pi)\in\mathbb{B}_{R}^{N}\times\Delta_{N}}{\sup}\left|\operatorname{F}_{\mathbb{D}}\left(\mu_{n}^{1},\ldots,\mu_{n}^{L},\sum_{i=1}^{N}\pi_{i}\delta_{Y_{i}}\right)-\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)\right|\\ &\leq\underset{(Y,\pi)\in\mathbb{B}_{R}\times\Delta_{N}}{\sup}2\left|\operatorname{F}_{W_{\epsilon,p}^{p}}\left(\mu_{n}^{1},\ldots,\mu_{n}^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)-\operatorname{F}_{W_{\epsilon,p}^{p}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)\right|\\ &+\frac{1}{L}\sum_{\ell=1}^{L}\left|W_{\epsilon,p}^{p}(\mu_{n}^{\ell},\mu_{n}^{\ell})-W_{\epsilon,p}^{p}(\mu^{\ell},\mu^{\ell})\right|.\end{split}$
In the right-hand side of the inequality, we bound the first term by Theorem 4.1 while the second term can be bounded thanks to the proof of (genevay2019sample, Theorem 3). In fact, their proof can be adapted almost directly to show that
$\mathbb{E}|W_{\epsilon,p}^{p}(\mu_{n},\mu_{n})-W_{\epsilon,p}^{p}(\mu,\mu)|=O\left(\frac{e^{\frac{{\mathcal{K}}}{\epsilon}}}{\sqrt{n}}\left(1+\frac{1}{\epsilon^{\lfloor d/2\rfloor}}\right)\right).$
This result is given for arbitrary measures $\alpha,\beta\in\mathcal{M}_{1}(\mathbb{R}^{d})$, their empirical counterparts $\hat{\alpha}_{n},\>\hat{\beta}_{n}$ and the quantity $W_{\epsilon,p}^{p}(\hat{\alpha}_{n},\hat{\beta}_{n})$, but at no point in the proof the independence of the samples upon which $\hat{\alpha}_{n}$ and $\hat{\beta}_{n}$ are constructed is mandatory. As a result, one can obtain the same bound with $\beta=\alpha=\mu$.
Appendix D Additional Lemmas
We provide here two well-known technical lemmas that are used throughout the main proofs in Appendix C. We then recall a classical result in optimal quantization.
D.1 Technical lemmas
Lemma D.1.
For $R>0$ and $p\geq 1$, we have
  • (i)
    the function $f:x\in\mathbb{B}_{R}\mapsto\|x\|^{p}$ is $pR^{p-1}$ Lipschitz,
  • (ii)
    the function $c:(x,y)\in\mathbb{B}_{R}\times\mathbb{B}_{R}\mapsto\|x-y\|^{p}$ is $\sqrt{2}p(2R)^{p-1}$ Lipschitz.
Proof.
(i) The case $p=1$ is just the result of triangular inequality. Second, by mean value theorem and the chain rule, one can show that for $p>1$,
$\nabla f(x)=p\|x\|^{p-2}x$
whose norm is upper bounded by $pR^{p-1}$.
(ii) Let $x,y,\in\mathbb{B}_{R}$. We first consider the case $p=1$ :
$\begin{split}\mid\|x-y\|-\|x^{\prime}-y^{\prime}\|\mid^{2}&=\mid\|x-y\|-\|x-y^{\prime}\|-(\|x^{\prime}-y^{\prime}\|-\|x-y^{\prime}\|)\mid^{2}\\ &\leq\mid\|x-y\|-\|x-y^{\prime}\|\mid^{2}+\mid\|x-y^{\prime}\|-\|x^{\prime}-y^{\prime}\|\mid^{2}\\ &\leq\|x-x^{\prime}\|^{2}+\|y-y^{\prime}\|^{2}\end{split}$
by triangular inequality.
Fore the case $p>1$, we differentiate the functional $c:(x,y)\mapsto\|x-y\|^{p}$, that is
$\begin{split}\partial_{x}c(x,y)&=\frac{p}{2}(\|x-y\|^{2})^{\frac{p}{2}-1}\partial_{x}\|x-y\|^{2}\\ &=p\|x-y\|^{p-2}(x-y).\end{split}$
It results that
$\|\partial_{x}c(x,y)\|=p\|x-y\|^{p-1}$
and therefore
$\begin{split}\|\nabla c(x,y)\|^{2}&=\|\partial_{x}c(x,y)\|^{2}+\|\partial_{y}c(x,y)\|^{2}\\ &=2p^{2}\|x-y\|^{2p-2}.\end{split}$
We then conclude by mean value theorem and the fact that $x,y,\in\mathbb{B}_{R}$.
Lemma D.2.
For any functions $f,g:\mathbb{R}^{d}\longrightarrow\mathbb{R}$ and subset $U$ of $\mathbb{R}^{d}$
$\left|\sup_{x\in U}f(x)-\sup_{x\in U}g(x)\right|\leq\sup_{x\in U}|f(x)-g(x)|$
and
$\left|\inf_{x\in U}f(x)-\inf_{x\in U}g(x)\right|\leq\sup_{x\in U}|f(x)-g(x)|$
Proof.
Let $x\in U$, then
$f(x)-\sup_{t\in U}g(t)\leq f(x)-g(x)\leq\sup_{x\in U}|f(x)-g(x)|.$
This being true for all $x\in U$ means that
$\sup_{x\in U}f(x)-\sup_{t\in U}g(x)\leq\sup_{x\in U}|f(x)-g(x)|.$
By inverting the role of $f$ and $g$ we can then show the first result. As for the second result for all $\overline{x}\in U$ we have
$\begin{split}\sup_{x\in U}|f(x)-g(x)|&\geq f(\overline{x})-g(\overline{x})\\ &\geq\inf_{x\in U}f(x)-g(\overline{x})\end{split}$
This being true for all $\overline{x}\in U$ means that
$\sup_{x\in U}|f(x)-g(x)|\geq\inf_{x\in U}f(x)-\inf_{x\in U}g(x).$
By inverting the role of $f$ and $g$ we can show the second result.
D.2 Lower bound Lemmas
In this subsection we prove that the worst-case lower bound given in bartlett2002minimax also holds for Wasserstein sparse barycenters.
Lemma D.3.
Let $\mathbb{D}=W_{2}^{2}$, the number of measures $L\in\mathbb{N}^{*}$ and the dimension $d\in\mathbb{N}^{*}$. Assume $N\geq 3$ and $n\geq 3435N$ and let $\sum_{i=1}^{N}\pi_{i}^{n}\delta_{y_{i}^{n}}\in\mathcal{M}_{1}^{N}(\mathbb{R}^{d})$ be a random measure depending on $nL$ random variables in $\mathbb{B}_{R}$. Then there exists $\mu^{1},\ldots,\mu^{L}\in\mathcal{M}_{1}(\mathbb{R}^{d})$ such that the estimation error of Corollary 4.4 verifies
$\mathbb{E}\left[\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}^{n}\delta_{y_{i}^{n}}\right)-\min_{(Y,\pi)\in\mathbb{B}_{R}^{N}\times\Delta_{N}}\operatorname{F}_{\mathbb{D}}\left(\mu^{1},\ldots,\mu^{L},\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)\right]\geq C_{d}\sqrt{\frac{N^{1-\frac{4}{d}}}{n}}$ (D.1)
where $C_{d}$ is an explicit constant which only depends on $d$.
Proof.
Let us first consider the case $L=1$. By (bartlett2002minimax, Theorem 1), there exists a measure $\mu\in\mathcal{M}_{1}(\mathbb{R}^{d})$ and an optimal quantizer $(Y^{*},\pi^{*})\in\underset{(Y,\pi)\in(\mathbb{R}^{d})^{N}\times\Delta_{N}}{\operatorname{argmin}}\ W_{2}^{2}\left(\mu,\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)$ such that
$\mathbb{E}\!\left[\min_{\pi\in\Delta_{N}}W_{2}^{2}\!\left(\mu,\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}^{n}}\right)-W_{2}^{2}\!\left(\mu,\sum_{i=1}^{N}\pi_{i}^{*}\delta_{y_{i}^{*}}\right)\right]\gtrsim\sqrt{\frac{N^{1-\frac{4}{d}}}{n}}.$
Since almost surely
$\min_{\pi\in\Delta_{N}}W_{2}^{2}\!\left(\mu,\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}^{n}}\right)\leq W_{2}^{2}\!\left(\mu,\sum_{i=1}^{N}\pi_{i}^{n}\delta_{y_{i}^{n}}\right),$
we deduce that
$\mathbb{E}\!\left[W_{2}^{2}\!\left(\mu,\sum_{i=1}^{N}\pi_{i}^{n}\delta_{y_{i}^{n}}\right)-W_{2}^{2}\!\left(\mu,\sum_{i=1}^{N}\pi_{i}^{*}\delta_{y_{i}^{*}}\right)\right]\gtrsim\sqrt{\frac{N^{1-\frac{4}{d}}}{n}}.$
The result for $L>1$ in (D.1) then follows by choosing $\mu^{1}=\cdots=\mu^{L}=\mu$. ∎
D.3 Optimal quantization
For completeness, we recall and prove this well known fact about optimal quantization.
Lemma D.4.
Let $\mu$ be an arbitrary measure and $Y\in(\mathbb{R}^{d})^{N}\setminus D_{N},$ then for all $p\geq 1$
$\underset{\pi\in\Delta_{N}}{\min}W_{p}^{p}\left(\mu,\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)=\int_{\mathbb{R}^{d}}\min_{i=1,\ldots,N}\|x-y_{i}\|^{p}d\mu(x).$
Proof.
Consider the functional
$(w,\pi)\in\mathbb{R}^{N}\times\Delta_{N}\longmapsto\int_{\mathbb{R}^{d}}\min_{i=1,\ldots,N}\{\|x-y_{i}\|^{p}-w_{i}\}d\mu(x)+\sum_{i=1}^{N}\pi_{i}w_{i}.$
It is concave in $w$ and convex in $\pi$, and we can therefore swap the min and the max in the following problem
$\min_{\pi\in\Delta_{N}}\max_{w\in\mathbb{R}^{N}}\int_{\mathbb{R}^{d}}\min_{i=1,\ldots,N}\{\|x-y_{i}\|^{p}-w_{i}\}d\mu(x)+\sum_{i=1}^{N}\pi_{i}w_{i}.$ (D.2)
Notice that a first order condition with respect to $\pi$ implies $w^{*}=0_{\mathbb{R}^{N}}$ at optimality. Let $\pi^{*}\in\Delta_{N}$ be a minimizer of (D.2). In order to use the duality expression (A.2), we consider $\tilde{\pi}\in\text{int}(\Delta_{M})$ the vector of weights $\pi^{*}$ whose null components have been removed. Following this construction and with a slight abuse of notation, we reorder the point cloud $Y:=(y_{1},\ldots,y_{N})$ so that the components $y_{i}$ corresponding to null weights have indexes $i\in[\![M+1,N]\!]$.
$\begin{split}\underset{\pi\in\Delta_{N}}{\min}W_{p}^{p}\left(\mu,\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)&=W_{p}^{p}\left(\mu,\sum_{i=1}^{M}\tilde{\pi}_{i}\delta_{y_{i}}\right)\\ &=\min_{\pi\in\Delta_{M}}\max_{w\in\mathbb{R}^{M}}\int_{\mathbb{R}^{d}}\min_{i=1,\ldots,M}\{\|x-y_{i}\|^{p}-w_{i}\}d\mu(x)+\sum_{i=1}^{M}\pi_{i}w_{i}\\ &=\max_{w\in\mathbb{R}^{M}}\min_{\pi\in\Delta_{M}}\int_{\mathbb{R}^{d}}\min_{i=1,\ldots,M}\{\|x-y_{i}\|^{p}-w_{i}\}d\mu(x)+\sum_{i=1}^{M}\pi_{i}w_{i}\\ &=\int_{\mathbb{R}^{d}}\min_{i=1,\ldots,M}\|x-y_{i}\|^{p}d\mu(x).\end{split}$
We can then consider two cases. First, suppose that $\mu$ does not give mass to the Voronoï set $\{x\in\mathbb{R}^{d}\>|\>\|x-y_{i}\|^{p}\leq\|x-y_{j}\|^{p},\>j\in[\![1,N]\!]\}$ for all $i\in[\![M+1,N]\!]$; this may be the case for example if $\mu$ is discrete. Then we have
$\int_{\mathbb{R}^{d}}\min_{i=1,\ldots,M}\|x-y_{i}\|^{p}d\mu(x)=\int_{\mathbb{R}^{d}}\min_{i=1,\ldots,N}\|x-y_{i}\|^{p}d\mu(x).$
Second, suppose that there exists $i\in[\![M+1,N]\!]$ such that $\mu$ gives mass to a set $\{x\in\mathbb{R}^{d}\>|\>\|x-y_{i}\|^{p}\leq\|x-y_{j}\|^{p},\>j\in[\![1,N]\!]\}$. By contradiction, suppose that $M<N$. Because $Y\notin D_{N}$ we then have
$\begin{split}\int_{\mathbb{R}^{d}}\min_{i=1,\ldots,M}\|x-y_{i}\|^{p}d\mu(x)&>\int_{\mathbb{R}^{d}}\min_{i=1,\ldots,N}\|x-y_{i}\|^{p}d\mu(x)\\ &=\max_{w\in\mathbb{R}^{N}}\min_{\pi\in\Delta_{N}}\int_{\mathbb{R}^{d}}\min_{i=1,\ldots,N}\{\|x-y_{i}\|^{p}-w_{i}\}d\mu(x)+\sum_{i=1}^{N}\pi_{i}w_{i}\\ &=\underset{\pi\in\Delta_{N}}{\min}W_{p}^{p}\left(\mu,\sum_{i=1}^{N}\pi_{i}\delta_{y_{i}}\right)\end{split}$
which contradicts the optimality of the minimum in $\pi$. We thus get the result. ∎